A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

22 May 2024 Version 1.13

Title word cross-reference

$104.00 [271]. $110.00 [282]. $111.85 [273]. $120 [272]. $130.00 [271]. $92.63 [276]. _{2} [108]. g [32]. M [12]. p_{2} [521].

-formula [32]. -quantile [12].

0 [273]. 0-367-63386-8 [273].

1 [560, 500]. 11 [425, 417, 415, 418, 423, 420, 416, 427, 429, 428, 426, 419, 424]. 125.00 [559]. 19 [265, 257, 412, 405, 406, 425, 422, 430, 243, 417, 415, 520, 411, 418, 404, 423, 402, 309, 420, 414, 416, 346, 403, 427, 410, 429, 407, 235, 428, 408, 409, 426, 448, 379, 419, 413, 424, 318, 380, 305, 347, 337].

1934-7 [186]. 1st [270, 560, 556, 582, 280].

2 [275, 269, 271, 589, 501, 268]. 2020 [67]. 2021 [252, 412, 405, 251, 406, 425, 417, 415, 411, 418, 250, 404, 423, 402, 420, 416, 403, 427, 410, 429, 407, 428, 408, 409, 426, 419, 424]. 294pp [559]. 2nd [20].

3 [270]. 366 [279]. 368pp [583].

4 [280, 284].

5 [123, 512, 282].

6 [283, 276].

7 [272]. 732066 [167].

8 [273, 278].

9 [412, 405, 406, 274, 411, 404, 402, 403, 410, 407, 408, 409]. 978 [283, 275, 269, 276, 272, 123, 270, 274, 512, 271, 560, 589, 282, 280, 278, 284, 268]. 978-0-367-14557-6 [283]. 978-0-367-18383-7 [272]. 978-0-367-19149-8 [278]. 978-0-367-19902-9 [274]. 978-0-367-20143-2 [269]. 978-0-367-25538-1 [583]. 978-0-367-35798-6 [276]. 978-0-367-36547-9 [559]. 978-0-367-46938-2 [268]. 978-1-003-27921-1 [560]. 978-1-032-04174-2 [582]. 978-1-032-25775-4 [556]. 978-1-032-25948-2 [589]. 978-1-118-64461-4 [284]. 978-1-119-54993-2 [279]. 978-1-119-58304-2 [271]. 978-1-138-04637-5 [282]. 978-1-138-59963-5 [123]. 978-1-138-71533-2 [275]. 978-1-4398-8538-3 [270]. 978-1-5095-3602-3 [281]. 978-3-030-22627-5 [512]. 978-3-030-67583-7 [285]. 978-981-15-5032-4 [280]. 9788-3-030-67582-0 [285].

Aaditya [131]. Aaron [500]. abatement [108]. abundance [571, 482, 248]. academic [82]. accelerated [310]. access [237, 342]. account [518]. Accounting [312, 3]. Accurate [440]. across [11]. activities [515]. activity [257, 236, 566, 197]. activity-based [257]. actor [111]. acyclic [88]. Adam [423]. Adams [583]. Adaptable [293]. Adaptive [99]. added [62]. additive [552]. address [332, 331]. addressing [255, 76]. adherence [76]. adjusted [9, 61]. Adjusting [10, 370, 15]. adjustment [58, 73]. adjustments [36]. administered [299]. Administrative [439, 531, 538, 533, 536, 539, 534, 543, 530, 540, 339, 485, 537, 541, 532, 535, 542]. admission [78]. Adulthood [300]. adults [383]. Advanced [586, 221]. advantage [383]. Advertising [440, 242]. advice [208]. advice-seeking [208]. affect [302]. African [311]. after [47]. AG [285]. against [325, 39]. age [171]. age-period-cohort [171]. ageing [256, 394, 196]. aggregated [235]. aggregation [175, 158]. AI [96]. aid [165]. air [547, 86]. air- [547]. airport [379]. al [500, 501]. al. [499, 502, 495, 496, 498, 503, 504, 505, 506, 420, 410, 429, 497, 419, 507, 508]. Alaska [110]. Alcohol [452, 205]. Alcohol-Related [452]. Alessio [356]. Alex [281, 354]. Algorithm [499, 502, 495, 496, 498, 503, 504, 509, 494, 505, 506, 497, 500, 501, 507, 508]. Algorithm-Assisted [499, 502, 495, 496, 498, 503, 504, 509, 494, 505, 506, 497, 500, 501, 507, 508]. Algorithmic [454]. algorithms [276]. Allan [67]. Allin [538]. allometric [11]. also [15]. ambidextrous [591]. ambient [234, 86]. America [572]. among [567, 383]. analyse [544]. analyses [66]. Analysing [439, 229]. Analysis [208, 249, 476, 587, 590, 272, 492, 451, 98, 491, 516, 557, 556, 335, 23, 20, 553, 559, 586, 469, 453, 521, 32, 227, 4, 273, 276, 260, 344, 517, 211, 110, 49, 327, 203, 308, 29, 205, 112, 84, 219, 231, 60, 369, 519, 315, 317, 73, 10, 75, 81, 63, 279, 42, 565, 229, 83, 259, 173, 167, 274, 474, 222, 460, 456, 280]. Analytic [268]. Analytics [582, 558]. Analyzing [190, 275]. Anastasopoulis [274]. anchoring [105]. Andre [497]. Anna [353, 142]. annual [572]. ANOVA [195]. answer [301]. Answers [596, 297]. ante [379]. Anthony [390, 489]. Antonella [536]. appear [261]. Application [499, 450, 478, 419, 413, 242, 574, 338, 502, 91, 495, 496, 498, 171, 12, 503, 504, 509, 494, 505, 506, 16, 111, 38, 72, 18, 14, 256, 394, 497, 500, 501, 201, 318, 507, 508, 380, 305, 210, 374]. Applications [24, 555, 476, 587, 221, 470, 547, 227, 289, 283, 53, 193, 271, 292, 70, 284]. Applied [270, 586, 272, 271]. apprenticeship [228]. Approach [218, 282, 475, 453, 310, 571, 312, 574, 52, 552, 572, 13, 45, 247, 46, 65, 37, 112, 36, 114, 484, 236, 544, 5, 309, 570, 162, 235, 116, 66, 97]. approach-from [247]. approaches [202, 510, 74, 42]. approximate [115]. Apps [435]. Arashi [284]. Area [455, 450, 546, 65, 12, 545, 7]. area-level [545]. areas [34, 14]. Armstrong [275]. Arnold [282]. art [340]. Arthur [134]. Artificial [477]. Aspects [349, 352, 353, 354, 355, 356, 357, 358, 359, 360, 361, 362, 363, 351, 364, 348, 350, 365, 368, 366, 367, 99]. Assessing [526, 422, 82, 381, 239]. Assessment [555, 499, 454, 574, 502, 495, 496, 498, 503, 504, 505, 506, 72, 261, 497, 10, 500, 501, 507, 508, 509]. assessment* [494]. assignments [71]. Assisted [499, 502, 495, 496, 498, 503, 504, 509, 494, 505, 506, 497, 500, 501, 507, 508]. associated [335]. association [53]. associations [264]. asymmetry [36]. Asymptotic [259]. ATE [154]. attack [241]. attainment [40]. Attitude [596, 297]. attractors [523]. attributable [77]. Attributes [578]. attrition [164]. Audit [157]. auditing [152]. Author [412, 509, 411, 420, 410, 429, 151, 419]. Authoritie [419]. Authorities [420, 414, 413]. authors [101]. auto [6]. auto-regressive [6]. autoregressions [575]. autoregressive [346]. auxiliary [160, 342]. available [64]. average [228, 112, 212]. averaging [162].

B [269, 284]. B/W [269]. Baffour [533]. Baillie [488]. Bakker [275]. balance [526, 372, 78]. balanced [17]. Banjamin [278]. Barbara [130]. Barndorff [391]. Barndorff-Nielsen [391]. Barrie [319]. Barry [390]. barycenter [550]. Baseball [268]. Based [269, 447, 187, 453, 257, 293, 91, 524, 152, 45, 334, 37, 377, 112, 335, 303, 510, 10, 63, 196, 467, 478, 247]. Basic [492, 110]. Basketball [268]. Basu [358]. battle [49]. Baumer [278]. Baye [188]. Bayesian [97, 245, 419, 34, 571, 521, 376, 574, 88, 572, 90, 13, 577, 110, 46, 37, 480, 369, 261, 378, 75, 81, 379, 515, 42, 576, 413, 383, 452, 273, 219, 464]. be [174, 15]. Beauty [26]. Becoming [446]. bee [571]. Beginner [586]. Behaviour [442, 13]. behavioural [547]. behaviours [13]. Being [472, 163, 83]. Benchmarks [588]. bereavement [264]. Berman [594]. Bernadini [536]. Bernard [533]. beta [55, 546]. betting [138, 128, 129, 127, 139, 132, 136, 140, 141, 142, 143, 144, 145, 133, 130, 134, 131, 146, 151, 126, 137, 148, 135, 149, 150, 147]. between [87, 53, 565, 54]. Beyond [154, 273, 270]. Bias [451, 10, 164, 3]. biases [115, 10]. bid [306]. bid-rigging [306]. Big [96, 432, 580, 23, 558, 162, 323, 157, 336, 347, 116]. big-data [157]. bike [313]. bike-sharing [313]. bilateral [572]. Bill [386]. Billé [328]. Billions [122]. Bin [559]. Binbing [283]. biostatistician [489]. Biostatistics [283, 561]. bipartite [527]. birth [295]. Bithell [67]. Bivariate [586, 271, 279]. block [526]. Boca [269, 123, 270, 167, 589, 282, 277]. body [52]. Bolin [556]. Bond [434, 264]. Bonferroni [179]. Book [96, 24, 474, 471, 583, 95, 19, 68, 246, 283, 249, 97, 190, 275, 476, 269, 273, 276, 466, 468, 462, 272, 459, 473, 99, 247, 98, 124, 220, 123, 270, 188, 458, 221, 461, 167, 245, 465, 274, 472, 26, 222, 271, 470, 285, 219, 457, 189, 560, 556, 22, 281, 23, 20, 122, 121, 248, 218, 477, 464, 244, 589, 559, 467, 21, 460, 582, 25, 280, 277, 187, 223, 191, 278, 475, 478, 284, 279, 456, 463, 469, 268]. border [5]. borne [547]. bound [108]. boundaries [71]. Bounding [373]. Bracher [417]. Bradley [445]. brain [87]. Brand [76]. Brazilian [343, 55]. break [544]. Brewer [215]. Brexit [163]. British [419, 413]. Broemeling [273]. Bryan [282]. budget [481]. burden [480]. Bynner [592].

C [269, 402]. C.P [167]. calculation [325]. calibration [238]. Calyampudi [563]. Can [40]. cancer [64]. Cane [385]. car [308, 516]. carbon [481, 234]. Cardiovascular [87]. Care [122, 324, 78, 196]. Carroll [275]. cartels [306]. case [211, 55]. cases [265, 309]. catchability [547]. categorical [34]. Category [450]. Catholic [50]. Causal [89, 477, 85, 77, 87, 106, 88, 338, 170, 180, 82, 71, 72, 64, 570, 373, 70]. causation [91]. cause [229]. cause-specific [229]. Causes [90, 378, 40]. Cautionary [581, 174, 259]. celebration [289, 287, 288, 291, 290]. Celeux [167]. censored [203]. censoring [241]. Census [14, 572]. Chai [138, 352, 539]. chain [85]. Challenge [429]. Challenges [292, 421, 70, 193]. Cham [512]. Change [49, 574, 8]. change-point [574]. changepoint [201]. Changliang [365]. Chapman [583, 283, 275, 269, 276, 123, 167, 274, 560, 589, 282, 278]. Characteristics [596, 297, 182, 181]. Charge [566]. chat [137]. checks [3]. Chen [272, 285]. Child [288, 295, 378, 83]. childhood [205, 54]. childless [383]. children [565, 33]. China [265, 83]. Chind [534]. Chloe [140]. Choi [353, 354, 142]. Choice [458, 262, 275]. Chris [333]. Christian [136]. Christine [138, 352, 539]. Christopher [583, 275, 92, 418]. chronic [199, 480]. Chuang [365]. Chun [543]. churn [231]. Cibella [536]. Circulant [461]. circumstances [54]. cities [195]. citizen [47]. city [11]. Claim [24, 483]. class [52, 576]. classical [42, 175]. Classification [23, 121]. classifications [44]. classifier [370]. clean [108]. climate [549]. Clinical [99, 448]. Closer [261]. Closures [422]. cluster [18, 64]. cluster-correlated [18]. clustered [335, 568]. Clustering [486, 206, 524, 13, 121, 38, 303, 63]. CO [108]. Cobham [281]. coefficient [12, 36]. coefficients [115]. coevolving [523]. cognition [205]. cognitive [87, 170, 50]. cohort [46, 171, 295, 66]. Colin [591]. collapse [311]. Collection [596, 435, 294, 297, 292, 115, 300, 336]. collections [199]. college [82]. combined [256, 394]. Combining [176, 199, 10, 104]. comments [259]. Communication [22, 138, 129, 139, 132, 136, 140, 141, 142, 143, 144, 145, 133, 130, 134, 131, 146, 151, 126, 137, 148, 135, 149, 150, 147, 128, 127]. commuter [13]. Companion [220, 467]. compare [235]. Comparing [260]. Comparison [446, 11, 104, 175, 305, 173]. comparisons [258, 61, 459]. competency [375]. competition [548]. competitive [526]. compiled [531, 538, 533, 536, 539, 534, 543, 530, 540, 537, 541, 532, 535, 542]. component [259]. composite [258]. compositional [60, 229]. Comprehensive [556]. compulsory [82]. Computation [464]. Computational [218, 282]. computationally [201]. Computer [220, 26]. concentrations [234]. Concepts [581, 558]. Conceptual [429, 421]. concern [200]. conditional [311]. conditions [227]. conduct [112]. confession [194]. confidential [339]. confounders [15]. Confounding [559, 312, 58, 55]. Congestion [566]. connectedness [243]. Consent [441]. consequences [40]. considering [308]. Consistent [158]. consumer [207, 157]. Consumption [436, 576]. contagion [84]. contain [51]. container [117]. Content [596, 297]. Contents [102, 225, 396, 479]. context [202]. Contextual [451]. Conti [499]. continuous [204, 205, 568]. Contribution [405, 406, 425, 417, 415, 418, 404, 423, 402, 416, 403, 407, 428, 408, 409, 426, 424, 531, 538, 533, 536, 349, 138, 352, 539, 534, 353, 499, 354, 128, 502, 129, 495, 127, 496, 355, 498, 356, 357, 139, 132, 136, 540, 503, 504, 358, 505, 140, 141, 359, 360, 506, 142, 537, 361, 143, 144, 145, 362, 363, 133, 130, 427, 134, 541, 497, 131, 500, 146, 351, 532, 364, 535, 501, 348, 507, 148, 135, 149, 350, 150, 508, 542, 365, 147]. contributions [289, 287, 288, 291, 290]. Controlled [562, 84]. Cook [354]. Cookbook [223]. copula [317]. Corona [293]. coronavirus [544]. Correction [328, 596]. correlated [545, 18, 522]. Corrigendum [253]. cost [262, 219, 173, 81]. cost-effective [173]. cost-effectiveness [219, 81]. Costs [452]. Council [489]. Count [24, 200, 339, 483, 235]. counteracting [515]. counts [195, 203]. county [482]. county-level [482]. Course [245, 315, 206]. CoV [411, 400]. Covariate [210, 204, 78, 378]. covariates [521, 372, 51, 171, 16, 233]. COVID [273, 354, 420, 349, 352, 353, 355, 356, 357, 358, 359, 360, 361, 362, 363, 351, 364, 348, 419, 350, 365, 265, 257, 412, 405, 406, 422, 243, 520, 411, 418, 404, 402, 309, 420, 414, 346, 427, 410, 429, 407, 235, 428, 408, 409, 426, 448, 379, 419, 318, 380, 305, 347, 337, 425, 430, 417, 415, 368, 423, 416, 366, 403, 413, 424, 367]. COVID-19 [273, 354, 420, 349, 352, 353, 355, 356, 357, 358, 359, 360, 361, 362, 363, 351, 364, 348, 419, 350, 365, 265, 257, 412, 405, 406, 422, 243, 520, 411, 418, 404, 402, 309, 420, 414, 346, 427, 410, 429, 407, 235, 428, 408, 409, 426, 448, 379, 419, 318, 380, 305, 347, 337, 425, 430, 417, 415, 423, 416, 403, 413, 424, 368, 366, 367]. Cox [393]. CR [563]. Crane [129]. crashes [308, 516]. CRC [583, 283, 275, 269, 276, 123, 270, 167, 274, 560, 556, 589, 559, 282, 582, 280, 277, 278, 268]. Credit [376].

Crime [39, 316, 110, 560]. Criminal

[451, 454, 197]. Criminologist [560]. crisis [257]. CRONOS [296, 302]. Cross [440, 5, 373, 296]. cross-border [5]. Cross-National [440, 296]. crowding [262]. cultural [514]. culture [514]. Cultures [514]. cumulative [383]. cure [16, 283]. Currie [389]. Curve [173, 383]. curves [381]. Cyber [471].

D [273, 123, 132, 403, 277]. Dagnelie [388]. Dagpunar [405]. daily [31]. Daniel [271, 278, 279]. Danny [541]. Dark [189]. darts [35]. Data [96, 436, 443, 249, 97, 587, 590, 492, 493, 98, 557, 465, 440, 193, 222, 432, 580, 435, 439, 23, 554, 589, 447, 441, 460, 584, 581, 7, 187, 475, 586, 442, 463, 558, 268, 531, 538, 204, 533, 262, 521, 536, 294, 227, 316, 572, 539, 4, 190, 276, 534, 518, 517, 543, 530, 89, 37, 549, 199, 308, 516, 291, 180, 540, 274, 200, 480, 485, 189, 177, 335, 60, 575, 62, 537, 212, 484, 161, 292, 18, 5, 194, 115, 64, 345, 541, 568, 162, 235, 183, 10, 75, 81, 300, 63, 379, 191]. data [278, 279, 532, 535, 229, 576, 323, 318, 209, 569, 542, 196, 155, 157, 259, 336, 347, 156, 116, 123]. Data-Based [447, 10]. Data-Driven [463, 7, 5]. database [161]. databases [339, 235]. David [275, 393, 389, 216, 488, 487]. Davis [67]. Death [554, 378, 346]. deaths [49, 309, 39]. Decision [499, 255, 571, 502, 495, 496, 498, 503, 504, 509, 494, 505, 506, 497, 500, 515, 501, 507, 508, 173]. Decision-Making [499, 495, 496, 498, 503, 504, 509, 494, 505, 497, 500, 501, 507, 508]. decisions [33]. Decoded [561]. decomposition [236]. Deeper [194]. default [376]. Delaunay [517]. demand [31]. dengue [230]. Denis [271, 279]. Denmark [108]. Density [303, 11]. Density-based [303]. deoxygenation [234]. dependence [547, 242, 483]. dependent [18, 239, 323, 318]. Derek [267]. Derive [123]. Design [456, 344, 299, 172, 29, 56]. Designing [154]. designs [310, 204]. detailed [10]. detect [56]. Detection [269, 574, 544]. Determinants [446, 442, 32]. determination [152]. Development [295, 322, 210, 258, 457]. developments [193, 292, 70]. deviation [117]. diabetes [87, 199]. Diagnostic [445, 317]. Dicing [554]. Did [112]. dietary [576]. differences [43]. different [235, 164]. Differential [589]. differentially [183]. Diggle [355, 406, 425]. Digital [436, 442, 548, 53]. Dimensional [466, 169, 190, 201]. dimensions [54]. Din [272, 285]. Ding [272, 285, 498]. Ding-Geng [272, 285]. dioxide [234]. Direct [446, 73, 28]. directed [88]. directions [340]. Dirichlet [8]. Dirichlet-multinomial [8]. dis [383]. disaggregation [62, 382]. discontinuity [310]. Discovering [88]. Discovery [473, 89]. discrepancy [238]. discrete [53]. Discussion [531, 538, 30, 533, 536, 349, 352, 353, 412, 354, 405, 355, 406, 425, 430, 543, 356, 417, 357, 415, 358, 411, 418, 368, 404, 423, 359, 360, 402, 361, 362, 363, 420, 416, 366, 403, 427, 410, 429, 407, 541, 428, 408, 409, 426, 351, 532, 364, 535, 332, 419, 424, 350, 367, 365, 539, 534, 499, 502, 495, 496, 498, 540, 503, 504, 509, 505, 506, 537, 497, 500, 501, 507, 508, 542, 138, 128, 129, 127, 139, 132, 136, 368, 140, 141, 142, 143, 144, 145, 133, 366, 130, 134, 131, 146, 151, 137, 348, 148, 135, 149]. Discussion [150, 367, 147]. Disease [457, 547, 199, 480, 277]. diseases [273, 198]. Disentangling [523]. Disorders [452]. dispersion [483]. distance [206]. distances [514]. distillation [137]. distributed [183]. distribution [304]. distributional [80]. distributions [160, 117]. districts [39]. DNA [190]. Do [115, 566, 302, 449]. Does [316, 48]. Domain [209]. domains [178]. domestic [257, 373]. Dong [351]. Don’t [373, 189]. double [114]. downscaling [235]. downside [57]. dowry [39]. Dr [534]. Draper [392]. Driven [463, 5, 7]. drives [13]. Drug [457, 322]. due [302]. Dunne [531, 538, 533, 536, 539, 534, 543, 540, 537, 541, 532, 535, 542]. during [293, 544, 115, 137, 337]. dyadic [568]. Dynamic [304, 525, 446, 104, 548, 262, 571, 481, 227, 111, 79, 159, 568, 116, 241]. dynamically [527]. dynamics [237, 573, 300].

e-book [269, 271]. early [230]. earnings [485]. earthquake [47, 73]. East [314]. eBook [275, 285, 560, 556, 280]. Ecological [411, 400]. Econometric [234, 274, 81]. Economic [434, 447, 32, 574, 529, 37, 60, 484, 236, 566, 346, 383, 54]. Economy [438]. ed [556]. edge [522]. Edition [476, 560, 283, 582, 280]. editor [321, 489]. edn [270]. Education [446, 164, 56]. education-entertainment [56]. educational [290, 40]. Effect [422, 376, 227, 110, 170, 31, 112, 163, 212, 314, 73, 373, 154, 15, 3]. effective [573, 173]. effectiveness [219, 78, 81]. Effects [449, 106, 242, 107, 307, 90, 109, 228, 171, 180, 84, 86, 114, 184, 82, 71, 570, 256, 394, 181, 17, 15, 298]. Efficient [413, 201, 419]. Efrén [502]. Ehsanes [284]. Eiler [391]. Eleanor [537]. elicitation [175]. eligibility [170]. elite [118]. emerging [198]. emissions [234, 86]. Empire [311]. empirical [104]. Employment [446, 485]. Ended [596, 297]. Ends [449]. energy [108]. Engineering [220]. English [526, 171]. enhancement [338]. Enhancing [194]. Enough [440]. enrolment [74]. ensemble [230, 544, 104]. entertainment [56]. entropic [6]. Epidemic [410, 399, 419, 413, 381, 314]. Epidemics [412, 401]. epidemiologic [66, 569]. epidemiological [312]. epilepsy [489]. Equation [188, 54]. Equations [589].

Equity [585]. Error [443, 204, 516, 233]. errors [232, 485]. establishing [81]. Establishment [439, 298, 113]. estimands [77]. estimate [262, 572, 114, 378, 311, 207]. estimates [531, 538, 533, 536, 539, 534, 543, 530, 343, 540, 537, 541, 532, 535, 542]. Estimating [106, 307, 232, 153, 437, 342, 337, 482, 71, 5, 17, 569, 228, 233]. Estimation [547, 455, 450, 199, 180, 411, 400, 14, 256, 394, 421, 263, 242, 108, 546, 179, 65, 12, 573, 545, 62, 212, 345, 7, 284, 569, 15, 233, 429]. estimator [91]. estimators [77]. Ethics [581]. Eugenie [503]. EUR [280]. Europe [552]. European [11, 195, 5]. evaluating [174, 75]. Evaluation [499, 441, 246, 574, 502, 495, 496, 498, 503, 504, 509, 494, 505, 506, 497, 500, 501, 507, 508]. evaluations [37, 112, 74, 164]. event [310, 525, 53, 241, 153, 380]. event-rates [153]. events [182]. Everything [96]. Evidence [462, 447, 246, 1, 253, 48, 381, 10, 83, 316, 110, 295, 566, 302, 247]. Evidence-based [247]. evidential [247]. evolution [8]. ex [379]. examine [327, 383]. Examining [87]. Examples [249, 245, 582]. excellence [321]. exchange [544]. exchanges [565]. exciting [520]. Exercises [587]. exogenous [346]. Expectile [450]. expenditure [157]. experience [302]. experiment [301, 56]. Experimental [444, 494, 210, 70, 73, 280, 499, 502, 495, 496, 498, 503, 504, 509, 505, 506, 497, 500, 501, 507, 508]. expert [552, 51, 44]. expert-preferences [552]. explain [327]. Explanatory [327]. Exploiting [161]. Explore [327]. Exploring [589]. exponential [260, 206]. exponential-distance [206]. exponential-family [260]. exposure [376, 312, 205]. exposure-penalized [312]. exposures [312]. expression [190]. extensive [107]. external [370, 66]. extra [82]. extrapolation [33]. extreme [182, 118].

F [67]. Fabrizio [536]. face [211]. Facebook [440]. factor [317, 75, 116]. Factors [462]. failure [310]. family [260]. Farcomeni [356]. Feasibility [444]. Fertility [447, 32, 46, 16, 315]. fertilizer [227]. figures [337]. Filtering [200]. Financial [249, 124, 109]. Finding [50]. findings’ [30]. Fine [419, 413]. finite [345]. First [368, 366, 367, 283, 359, 362, 300, 347, 349, 352, 353, 354, 355, 356, 357, 358, 359, 360, 361, 362, 363, 351, 364, 350, 365]. Fitting [341]. fixed [184]. FL [269, 589, 282, 277]. flagging [306]. Flavia [120]. Flexible [80]. flows [572]. fluctuations [48]. follow [295]. follow-up [295]. following [208, 316]. food [255, 237]. Football [268]. Footprint [448]. Force [450, 343, 14, 337]. Forecasting [590, 434, 46, 108, 45, 577, 60, 346]. Forensic [451, 246, 238]. forensics [53]. Foreword [287]. Forks [454]. Formalisation [451]. forms [161]. formula [32]. forty [526]. Foundations [21]. founding [489]. fractal [48]. fraction [77]. Framework [443, 445, 4, 514, 173, 64, 33, 336]. France [447]. Frances [277]. Francis [67]. fraternal [264]. Fraud [122]. Fred [274, 333]. Freda [387]. freemium [231]. Freight [57]. frequency [6, 575]. Fresh [475]. Fricker [277]. friendly [339]. Frühwirth [167]. Fulfilling [47]. fuller [194]. Functional [195, 557, 65, 313, 229]. functioning [170]. functions [304, 12]. Fund [110]. Future [584, 529, 292].

G [167]. G. [319]. gained [313]. games [231]. gap [551, 65, 236]. gaps [156]. Garib [364]. gas [31]. Gasser [593]. Gaussian [88, 17]. Gavin [415]. GBP [123, 167]. GDP [162]. Ge [357, 361]. Gender [43, 551, 236]. gene [190, 89]. general [3]. generalised [346]. generalizable [28]. generalization [154]. Generalized [270]. Generalizing [1, 210, 253]. Generating [183]. generation [517]. generic [76]. genetics [40]. Geng [272, 285]. Gentleman [321]. geographical [34]. Geostatistical [557, 379]. Geraci [356]. Gerald [93]. Gerbing [123]. German [293, 527]. Germany [243, 447]. gestational [205]. Ghosh’s [502]. Gianpaolo [424]. Gibson [415]. Gilmore [214]. Glenn [138, 129, 139, 132, 136, 140, 141, 142, 143, 144, 145, 133, 130, 134, 131, 146, 137, 148, 135, 149, 150, 147, 128, 127]. global [237, 481, 111, 309, 201, 198]. Golam [284]. Goldstein [67, 289, 287, 288, 291, 290]. Goodhardt [93]. governance [165]. grant [174]. graph [88, 260]. graphs [85]. great [591]. Greenland [139]. grid [517]. Griffiths [427]. Group [510, 277, 169]. Group- [510]. grouped [203, 177, 209]. groups [60, 515]. groupwise [570]. growing [527]. Growth [410, 399, 383]. Grünwald [132]. Guide [560, 586, 44]. Guy [416].

H [67, 556]. Haigh [185]. hail [239]. Hall [583, 283, 275, 269, 276, 123, 167, 274, 560, 589, 282, 278]. Hall/CRC [283, 275, 269, 276, 123, 167, 274, 560, 589, 282, 278]. hand [35]. Handbook [562, 474, 476, 459, 124, 167, 582]. Handbooks [167]. handle [37]. Handling [493]. Hardback [283, 275, 560, 556, 273, 276]. Hardbound [269]. hardcopy [280]. Hardcover [589, 271, 285]. Hare [275]. Harmful [449, 83]. Harold [488]. Harry [129]. Harvey [67, 289, 287, 288, 291, 290]. Haslam [252]. having [233]. Hawkes [321]. hazard [211]. hbk [583, 281, 582]. healing [81]. Health [87, 165, 465, 86, 122, 321, 107, 37, 520, 324, 202, 510, 74, 280, 277, 3, 83, 54, 288]. healthcare [574, 314, 3]. heart [241]. Heartbeat [438]. heat [202, 182]. heat-health [202]. hedonic [9]. Heidegger [325]. helped [489]. Helsinki [313]. Hennig [136]. Henry [67, 540]. Heterogeneity [551, 447, 111, 261, 154]. Heterogeneous [412, 401, 547, 570, 305]. hidden [237, 71]. hierarchical [521, 46, 308, 105, 378, 383]. High [466, 208, 169, 316, 190, 344, 50, 201]. High-Dimensional [466, 169, 190, 201]. high-influence [344]. high-profile [316]. Hills [186]. historical [49, 153]. History [25, 241, 63]. HIV [567, 437]. Hoboken [271, 284]. Hobokon [279]. holidays [31]. Holt [487]. Horten [278]. Hospital [452, 521, 18]. hot [35]. hotels [548]. Housing [444]. Howard [275]. Human [499, 502, 495, 496, 498, 503, 504, 509, 494, 505, 506, 497, 500, 501, 507, 508]. human-decision [502]. human-decision-making [506]. Humanity [277]. Hunsicker [503]. hurdle [55]. Hutton [504]. hyperevent [528]. hypergamy [373].

Iain [389]. Ian [428]. identical [264]. Identification [84, 484, 340]. identify [173, 202]. Identifying [31, 576, 52]. II [275]. illustration [4]. Illustrations [269]. Imagery [433]. images [301]. Imai [499, 502, 495, 496, 498, 503, 504, 505, 506, 497, 500, 501, 507, 508]. Impact [412, 401, 108, 295]. impacts [520]. impairment [87]. Imperfect [21]. implementation [208, 283]. implications [71]. Improved [4, 231]. Improving [66]. Imputation [492, 371, 176]. in-person [294]. incentive [302]. Incentives [447, 566]. incidence [201]. Inclusion [552, 16]. Income [300, 511, 110, 209, 342]. Incomplete [493, 89, 371, 480]. Inconsistency [454]. incorporate [116]. Incorporating [573, 575]. increasing [331, 330, 332]. Index [101, 52, 179, 157]. India [484, 373, 39]. Indian [544]. Indicators [455, 342]. indices [346, 158].

Indirect [446, 449]. Individual [447, 312, 307, 184, 261, 510]. individual-based [510]. individual-level [312, 261]. individuals [52]. Indonesia [379]. induced [15]. Inequality [177, 546, 179, 377, 510]. infant [165]. infectious [273]. Infer [448]. Inference [473, 21, 25, 419, 338, 240, 72, 70, 64, 85, 570, 191, 413, 157, 77, 477]. Inferring [38, 570]. influence [169, 344, 369]. Info [21]. Info-metrics [21]. inform [33]. Information [2, 27, 41, 69, 103, 125, 168, 192, 226, 254, 286, 329, 397, 431, 36, 160, 47, 199, 71, 342, 21]. Informative [410, 399, 113, 335, 341]. informativeness [338]. infrastructure [11]. infrastructures [293]. innovate [299]. Innovating [297, 596]. Innovative [457]. Inpatient [452]. Insights [313, 442, 440]. Instantaneous [419, 413]. Instead [444]. Institute [288]. instrument [50]. Instrumental [79, 80, 83]. instruments [78]. Insurance [588, 24, 471]. insurers [239]. integrated [482, 378]. integration [510]. Intelligence [477]. intensity [200]. intensive [107, 78]. inter [174]. inter-rater [174]. interaction [528]. interactions [523]. Interactive [187]. Intercensal [433]. interference [117]. Internet [293, 257, 5]. interplay [243]. interrelationships [54]. intervention [75]. Interventions [412, 401, 79]. Interviewer [109, 181]. Introducing [298]. Introduction [96, 587, 512, 247, 398, 276, 466]. Introductory [467]. Invention [584]. inventors [527]. inverse [1, 253]. Investigation [444]. investigations [64]. Investing [585]. Invited [415, 404, 423, 402, 416, 403, 424]. ISBN [583, 283, 275, 269, 273, 276, 272, 123, 270, 167, 274, 512, 271, 285, 560, 556, 281, 589, 559, 282, 582, 280, 278, 284, 279, 268]. ISBN-13 [512]. Issue [2, 27, 41, 69, 103, 125, 168, 192, 226, 254, 286, 329, 397, 431, 333]. issues [322]. Italian [434, 211]. item [160, 181]. ix [268].

J [355, 406, 425, 415, 271, 278, 279]. Jacob [560]. Jails [437]. Jana [358]. Jane [504]. January [186]. Japanese [374]. Jasmina [427]. Jennet [542]. Jenny [285]. Jewell [418]. Jiang [368]. Jocelyn [556]. John [67, 92, 390, 271, 589, 284, 592, 252, 405, 543, 185, 404]. Johnson [489]. joint [317]. jointly [231]. Jolliffe [120]. Jones [251]. Jorge [145, 362, 363]. Jose [505]. Joseph [93]. Journalism [22]. journey [291]. Jr [277]. judgment [275]. Judith [147]. Julie [270]. jumps [200]. June [412, 405, 186, 406, 425, 417, 415, 411, 418, 404, 423, 402, 420, 416, 403, 427, 410, 429, 407, 428, 408, 409, 426, 419, 424]. Justice [451, 454]. Justify [449]. Justin [402]. Jyoti [269].

Kane [282]. Kaplan [560, 278]. Karl [272]. Karlaftis [274]. Katie [540]. Kaushik [358]. Keiding [320]. Keith [275, 582]. Kemp [387]. Ken [215]. Kenneth [215, 140]. Kenya [91]. kernel [66]. key [567]. Khan [280]. Kibria [284]. kidney [199]. Kingdom [420, 255, 414, 207]. Kingman [404]. Kingsuk [358]. Knives [454]. Know [96, 189]. Knowing [28]. Knowledge [473, 188]. Korea [314]. Krakauer [140]. Kucharski [423]. Kuldeep [141, 359, 360, 506]. Kumar [141, 359, 360, 506]. Kwok [354].

L [353]. Labour [450, 343, 107, 14, 337]. Lagnajita [358]. Lai [353, 142]. Lan [269]. landmarking [104]. Language [436, 325]. large [262, 210, 527, 339, 575]. large-scale [262, 210]. Lars [217]. Latent [518, 525, 52, 260, 486, 576, 54, 523]. lattice [308, 516]. law [265, 117, 374, 537]. Lawrence [594]. League [526]. learned [294]. Learning [473, 439, 449, 469, 306, 114, 82, 202, 159, 104, 323, 380, 282, 583, 23, 218, 464, 244, 478]. Leeds [516]. left [241]. Legler [270]. lens [577]. Leonard [489]. lessons [294, 314]. Leung [353]. level [312, 316, 377, 545, 482, 335, 71, 261, 156]. Levelers [20]. leverage [36]. Leveraging [160]. Lewis [282, 166]. Lewnard [411]. Li [559, 543, 357, 361, 133]. Li-Chun [543]. Life [315, 206, 63, 54]. Life-course [315, 206]. lifespan [264]. lifetime [289, 287, 288, 291, 290, 156]. Likelihood [269, 247, 238, 241]. limitations [161]. limited [342]. line [376, 373]. Linear [270, 372, 171, 155, 470]. Linearization [179]. liner [117]. Linguistic [596, 297]. Lingwood [591]. Link [441, 550]. Linkage [437, 155, 291]. Linkage-data [155]. linked [4, 485, 207]. Linking [442]. Linton [357, 361]. Listening [438]. listings [161]. literacy [109]. literary [8]. live [137]. Living [554]. loans [38]. Local [420, 414, 419, 413, 170]. log [377]. log-normal [377]. logistic [260]. London [566]. Long [182]. Long-term [182]. Longevity [60, 156]. Longford [143]. Longitudinal [565, 208, 37, 295, 519, 206, 568]. Lorenzo [407]. loss [159]. Loynes [384]. Lyberg [217]. Lyle [273].

M [67, 266, 284]. Machine [439, 244, 202, 478, 469, 306, 114, 104, 380, 473]. macroeconomic [45]. mail [198]. Making [499, 502, 495, 496, 498, 503, 504, 509, 494, 505, 506, 497, 500, 501, 507, 508, 196]. Mallows [591]. Management [124]. manager [346]. Mannering [274]. Mannheim [293]. many [372]. Maori [233]. Mapping [379, 419, 413, 237, 377]. Marco [356]. marginal [160, 228, 15]. margins [107]. Markdown [223]. marked [239]. Market [434, 548]. markets [57]. Markov [237]. Martens [512]. Martin [144].

Martingale [468]. Mass [22, 52, 176]. matches [577]. matching [242, 74]. material [211]. Mateu [145, 362, 363]. Mathematical [55]. Mathematics [26, 268]. matrices [378, 461]. Mats [501]. matters [189]. Matthew [274]. Maurice [595]. maximum [241, 375]. may [174, 15]. McLennan [386]. McNulty [582]. Md [284]. Meaning [123]. Means [449, 336]. Measure [411, 257, 520]. Measurement [516, 578, 204, 109, 485, 177, 159, 233]. Measures [588, 400, 546, 258]. Measuring [511, 552, 108, 484, 340, 514, 248]. Media [436, 200, 523, 472]. median [65]. mediating [87]. Mediation [559, 32]. Medical [269, 20, 219, 375, 489, 244]. Medicine [489]. Meeting [349, 352, 353, 412, 354, 405, 355, 406, 425, 430, 356, 417, 357, 415, 358, 411, 418, 404, 423, 359, 360, 402, 361, 362, 363, 420, 416, 403, 427, 410, 429, 407, 428, 408, 409, 426, 351, 364, 419, 424, 350, 365, 30, 368, 366, 367]. Meets [432]. membership [197]. Memories [289, 290]. memory [333]. Meng [133]. Mervyn [119]. Meta [272, 317, 61, 10, 280]. Meta-Analysis [272, 317, 10, 280]. meta-regression [61]. Metcalf [402]. Metered [443]. method [230, 175, 15]. Methodologies [473]. Methodology [104, 269]. Methods [555, 283, 468, 493, 167, 562, 457, 580, 433, 559, 478, 268, 289, 274, 177, 241, 544, 79, 280]. methylation [190]. metrics [59, 21]. Michael [384, 252, 490, 282, 186]. Michel [512]. micro [156]. micro-level [156]. microdata [113, 340]. Microeconometrics [583]. mid [54]. mid-life [54]. Middle [314]. migration [572]. Mike [252]. minding [194]. mining [198]. minority [569]. misclassification [378, 370]. Misinterpretations [411, 400]. Missing [492, 164, 37, 567, 233]. misuse [482]. mitigations [346]. mixed [376, 298, 295, 299, 6, 575, 317, 300, 197]. mixed-frequency [575]. mixed-mode [298, 295, 299, 300]. Mixture [455, 167, 486, 377]. mixtures [304, 546, 206]. mobility [178, 243, 484]. modal [318]. mode [298, 295, 299, 300]. Model [24, 455, 524, 491, 420, 414, 63, 445, 453, 525, 34, 262, 237, 526, 481, 242, 232, 90, 327, 549, 377, 545, 482, 483, 6, 111, 159, 378, 317, 568, 313, 162, 75, 116, 379, 576, 383, 305, 54, 523]. Model-Based [453, 524, 63]. Modeling [473, 464, 589, 21, 582, 452, 285]. Modelling [32, 171, 527, 375, 558, 169, 304, 376, 234, 195, 178, 52, 289, 47, 12, 480, 485, 231, 567, 9, 345, 81, 182, 311, 305, 97]. Models [471, 412, 401, 270, 470, 244, 578, 310, 160, 521, 283, 307, 88, 275, 260, 518, 486, 51, 327, 516, 343, 105, 339, 335, 184, 16, 212, 241, 528, 206, 346, 55, 116, 341, 197, 77]. models-estimands [77]. Moderation [559, 106]. moderators [106]. Modern [167, 263, 42, 99, 278]. Modified [203]. modifiers [15]. Mohammad [284]. Monitoring [277]. month [343]. monthly [62, 162, 337]. Morgan [592]. Mortality [448, 304, 165, 256, 394, 229, 374]. motivation [231, 28]. moves [207]. Multi [450]. Multi-Category [450]. multidimensional [524, 305]. multierror [232]. Multilevel [178, 270, 519, 61, 289, 212, 55, 341, 54]. multinomial [51, 371, 8]. Multiple [459, 492, 270, 42, 233, 263, 199, 317, 75, 201, 54]. Multiple-systems [42]. multiply [518]. multisite [59]. multistate [480, 77]. multitrait [232]. multitrait-multierror [232]. multivalued [73]. Multivariate [24, 455, 308, 483, 567, 586, 481, 545, 271, 75, 279, 341, 383]. musical [114].

Nason [416, 366]. Nath [364]. National [440, 295, 296]. Natural [436, 372, 271]. Nature [285, 280, 48]. near [47]. Nearest [371]. Need [96, 47, 196]. negative [242, 81, 523]. neighbour [371]. neonatal [378]. Net [512]. Netherlands [178, 337]. Network [520, 571, 521, 260, 518, 527, 308, 516, 111, 346, 61, 379, 523]. Networks [522, 513, 208, 548, 524, 514, 517, 486, 519, 528, 303, 85, 550]. Neural [434]. News [436, 434]. Next [277]. Nicholas [278]. Nick [143]. Nicoletta [536]. Niels [320]. Nielsen [391]. Nigel [321]. NJ [271, 279]. noisy [518, 62, 484]. Non [435, 439, 106, 307, 552, 91, 171, 549, 76, 177, 176, 70, 115, 73]. non-additive [552]. non-adherence [76]. non-experimental [70, 73]. non-linear [171]. non-overlap [76]. non-parametric [307, 91, 549, 177]. Non-Participation [435]. non-probability [176]. non-randomized [106]. Non-Response [439, 115]. nonignorable [160]. Nonlinear [318]. nonparametric [576]. nonresponse [160, 298, 181]. normal [377]. Norman [392]. North [266]. note [174]. novel [52]. Nowcasting [162, 419, 413, 6, 575]. Number [411, 400, 584, 437, 573, 325]. Numbers [560, 410, 399, 429, 421, 584, 419, 413]. nursing [324].

obesity [171]. Object [460].

Observational [462, 453, 78, 75, 81].

obtain [372]. ocean [234]. O’Farrell [540]. official [116]. Ohio [482]. Ole [391]. Oliver [357, 361]. Olympics [55]. one [317, 83]. one-child [83]. one-factor [317]. O’Neill [403]. Online [440, 296, 293, 292]. only [531, 538, 533, 536, 539, 534, 543, 530, 540, 537, 541, 532, 535, 542, 83]. Open [596, 297, 81]. Open-Ended [596, 297]. operations [316]. opinion [51]. opioid [482]. opportunities [336]. opposing [13]. Optimal [74, 173]. Optimally [17]. Optimising [380]. Oral [596, 297]. order [260]. ordered [105]. ordinal [568, 380]. Oriented [460]. origins [311]. Osimani [130]. other [411, 400, 118, 57, 158, 268]. outbreak [230, 314, 198]. outbreaks [277]. Outcome [18, 204, 37, 371]. Outcome-dependent [18]. Outcomes [450, 310, 50, 38, 75, 380]. output [62]. overcome [241]. overlap [76]. overlapping [34, 62]. own [57]. Oyo [311].

P [583, 138, 352, 539]. Pages [275, 272, 283, 273, 276, 589, 280]. Pairwise [258, 305]. Panagiotis [274]. pandemic [257, 293, 544, 337, 349, 352, 353, 354, 355, 356, 357, 358, 359, 360, 361, 362, 363, 351, 364, 348, 350, 365, 368, 366, 367]. Panel [300, 294, 227, 293, 549, 184, 296, 302, 293]. Panovska [427]. Panovska-Griffiths [427]. Papers [405, 406, 425, 417, 418, 404, 423, 402, 416, 403, 427, 407, 428, 408, 409, 426, 424, 415, 70]. parameters [47, 483, 345]. Parametric [12, 453, 32, 307, 91, 37, 549, 177, 177]. parents [565, 383]. Parry [251]. Parry-Jones [251]. part [500, 501]. partial [484]. Partially [212]. participating [302]. Participation [450, 435]. partisanship [523]. partnership [54]. patents [527]. pathology [87]. patterns [576, 39]. Paul [270, 538, 535, 148]. pay [236]. pbk [281]. Peace [272]. Pedersen [134]. pedestrian [195]. peer [174]. Pellis [407]. penalized [312]. Peng [283, 498]. Pension [170]. People [582, 228, 199]. perceived [305]. perception [43]. perceptions [305]. Performance [72, 260, 71]. performances [82]. period [171]. Permanent [110]. person [294]. Personalised [196]. personality [315]. Persons [437]. perspective [261, 315]. Perspectives [585]. Pessimistic [412, 401]. Peter [355, 406, 425, 132, 250]. Pfeffermann [541]. phase [204]. Philip [266, 403]. Philogenetic [25]. Philosophy [188]. physical [566]. Pictures [444]. Pier [499]. Pierre [388]. placebo [84, 56]. placebo-controlled [84]. Plague [277]. planners [552]. plant [86]. player [231]. playing [231]. plotly [187]. point [574, 49, 520, 8, 239]. Poisson [203, 183]. Poisson-distributed [183]. Policy [222, 552, 74, 83]. Political [436, 596, 297]. politics [325]. Polity [281]. pollinator [571]. polyadic [528]. Poole [275]. pooled [212]. population [531, 538, 533, 536, 547, 539, 534, 543, 530, 540, 537, 345, 541, 61, 532, 535, 542, 342, 233, 77]. population-adjusted [61]. populations [567, 277]. positive [523]. Post [334, 73]. Post-strata [334]. post-traumatic [73]. Poverty [455, 377, 65]. Power [265, 517, 86, 380]. pp [269, 123, 270, 167, 274, 512, 271, 285, 560, 556, 281, 582, 284, 268, 277]. Practical [99, 324, 22]. Practice [492, 580, 114, 247]. Pradesh [39]. Pre [228]. Pre-apprenticeship [228]. precision [380]. Predicted [449]. Predicting [401, 184, 47, 567, 318, 412]. Prediction [244, 34, 276, 230, 550, 104, 209, 196, 473]. Predictive [454, 558, 169, 327]. Predictors [446]. predominantly [294]. Preface [70, 333]. preferences [552]. preferential [9]. Premier [526]. presence [180]. Presenting [443]. Preserving [433]. Presidential [331, 332]. Press [583, 269, 123, 270, 274, 556, 281, 589, 559, 582, 280, 277, 278, 268]. pressure [81]. presumed [164]. Pretrial [499, 502, 495, 496, 498, 503, 504, 509, 494, 505, 506, 497, 500, 501, 507, 508]. prevalence [199, 482, 567]. previous [302]. Price [272, 280, 157, 158]. pricing [548, 57]. primer [326]. principal [259]. principle [247]. prior [175]. Privacy [463]. private [161, 183]. Priyantha [150, 508]. Probabilistic [451]. probabilities [311]. probability [1, 253, 344, 293, 91, 176, 117, 220, 458, 475]. probability-based [293]. probit [184]. procedure [90, 201]. process [239, 17]. processes [520, 522]. Processing [436, 336]. produced [50]. production [108]. Productivity [11]. professional [35]. Professor [490]. profile [316]. profiling [18]. program [210]. programs [372]. progress [375]. progression [231]. Project [467]. Project-Based [467]. projection [91]. projection-based [91]. projects [574, 59]. ProMED [198]. ProMED-mail [198]. promote [566]. prompt [78]. propensities [115]. Propensity [205, 212, 17]. properties [118]. property [239]. proportions [14]. Proposal [331]. Proposer [531, 495, 127, 348]. Protecting [463]. Proxy [157]. Psychological [578, 83]. Psychometrics [579]. Public [499, 552, 502, 495, 496, 498, 31, 503, 504, 200, 509, 494, 505, 506, 43, 497, 500, 501, 507, 305, 55]. publications [194]. publicly [64]. purchases [5]. purchasing [346]. pure [569]. Python [464, 497, 586, 582].

Qingzhao [559]. quality [86, 194, 302]. quantification [263, 42]. Quantifying [257, 53, 346, 164, 156, 51, 309]. quantile [376, 12]. quantiles [334]. Quantitative [279, 33]. quarterly [62]. question [181]. Questions [596, 444, 301, 297]. Quirke [390].

R [587, 272, 473, 245, 512, 556, 22, 559, 467, 582, 187, 279, 354, 491, 583, 476, 276, 270, 271, 285, 560, 589, 278, 249, 190, 123, 221, 470, 223, 475, 469, 465]. R/Stan [491]. Radhakrishna [563]. rainfall [338]. Ram [269]. Ramdas [131]. random [260, 461]. randomised [380]. Randomized [99, 562, 106, 372, 1, 253, 84, 154]. Rank [478]. Rank-Based [478]. Ranking [118]. Rao [563]. Rare [457]. rate [343, 544]. rater [174]. Rates [410, 399, 567, 346, 153, 57, 374]. Ratio [269, 371]. ratios [238]. Raton [123, 270, 167, 589, 282, 277]. Rawlins [490]. re [340]. re-identification [340]. Real [577, 429, 421, 260]. Real-time [577]. real-world [260]. realistic [517]. realtime [47]. Reason [458]. Reconciling [485]. Record [437]. records [153]. reduce [316]. reducing [242]. Referees [100, 224, 564, 395]. referendum [163]. refined [78]. regional [243, 6, 575, 342]. regions [11]. Register [447, 196]. registers [207]. registries [569]. Regression [310, 476, 450, 270, 470, 582, 15, 376, 307, 12, 203, 285, 483, 61, 8, 80, 284, 318, 155, 556, 20]. Regression-with-residuals [15]. regressive [6]. regulation [89]. rejected [38]. Related [452]. Relational [528, 525]. relations [568]. relationship [87]. Relationships [411, 400]. relative [65]. reliability [263, 174]. Removing [169]. rental [161]. rented [161]. repeated [345]. replicating [30]. replication [172, 29, 59, 462]. Replies [430]. Reply [412, 411, 420, 410, 429, 419, 543, 509, 368, 366, 151, 367]. reported [518, 238]. reports [47, 485]. Representative [569]. Representativeness [296]. Reproduction [411, 400, 410, 399, 429, 421, 419, 413, 573]. Research [489, 440, 580, 435, 20, 290]. residential [207]. residuals [15]. respiratory [314]. respondent [181]. Response [439, 232, 105, 115, 346, 302]. responses [568]. responsive [299]. result [163]. Results [294, 300, 210, 10]. retention [231]. Rethinking [245]. retirement [170]. retrospect [347]. reveal [40]. Review [96, 24, 474, 471, 583, 246, 283, 249, 97, 190, 275, 476, 269, 273, 276, 466, 468, 462, 459, 473, 99, 247, 98, 124, 220, 123, 270, 188, 458, 221, 461, 167, 245, 465, 274, 472, 26, 222, 271, 470, 285, 219, 457, 189, 560, 556, 22, 281, 23, 20, 122, 121, 248, 218, 477, 464, 244, 589, 559, 282, 467, 21, 460, 582, 25, 280, 277, 187, 223, 191, 278, 475, 478, 284, 279, 456, 463, 469, 268, 307, 272, 174, 246]. reviews [95, 19, 68]. revolutionize [489]. Reynold [428]. Rice [140]. Richard [267, 392, 251]. Richardson [331, 408, 332]. ridge [284]. Rigdon [277]. rigging [306]. right [203]. right-censored [203]. Riley [409, 426]. Risk [471, 555, 124, 588, 454, 244, 152, 173, 241, 261, 256, 394, 239, 340, 569, 164]. risk-based [152]. risks [379, 57, 569]. risky [574]. Rita [595]. Roback [270]. Robert [167, 384]. Robinson [267]. Roger [213]. Rogna [328]. role [87, 231]. role-playing [231]. rolling [74]. Ronald [215, 277]. Rosenthal [275]. Rosina [385]. Roton [269]. rounding [372]. routine [199]. Royal [417, 415, 418, 423, 416, 427, 429, 412, 405, 406, 425, 411, 404, 402, 420, 403, 410, 407, 428, 408, 409, 426, 419, 424]. Royce [275]. RSS [430]. Rule [188]. rules [170, 44]. Ruodu [149]. rupee [544]. Ryan [275, 144].

S [167, 278, 594]. Safety [499, 502, 495, 496, 498, 503, 504, 509, 494, 505, 506, 43, 497, 500, 501, 507, 508]. Saleh [284]. Sample [344, 152, 34, 334, 370]. samples [176]. sampling [204, 1, 253, 344, 371, 113, 335, 18, 9, 341, 157, 456]. Sander [139]. sanitation [91]. Sarah [540]. SARS [411, 400]. SARS-CoV-2 [411, 400]. Sarvet’s [501]. SAS [559, 98]. Satellite [433]. saturated [339]. Save [122]. Saving [277]. Savvy [584]. scale [262, 210, 258, 58]. Scales [419, 413, 11, 235]. Schnakenburg [512]. Schnatter [167]. School [422]. schools [50, 55]. Schure [147]. Science [443, 440, 222, 432, 580, 581, 276, 47, 238, 193, 278, 279, 96, 220, 188, 465, 26, 25, 191]. Sciences [590, 556]. scientific [138, 128, 129, 127, 139, 132, 136, 140, 141, 142, 143, 144, 145, 133, 130, 134, 131, 146, 151, 126, 137, 148, 135, 149, 150, 147]. Scientists [269, 326, 271, 246]. score [205, 212]. scores [17, 512]. Scraping [437]. scratch [326]. screen [173]. screens [306]. search [257, 316]. seasonal [549, 159]. seasonality [256, 394]. seasons [526].

Sebastian [417]. Second [476, 360, 363, 370]. secondary [4]. Seconder [349, 128, 496, 532]. sector [161]. secular [76]. Secure [336]. security [255, 471]. seeking [208]. segments [57]. Seheult [67]. SEIR [412, 401]. seismic [211, 47]. Seismonomics [438]. Selecting [58]. Selection [25, 478, 90, 210, 378, 3]. self [520, 299, 3]. self-administered [299]. self-exciting [520]. self-selection [3]. Semi [453, 205]. semi-continuous [205]. Semi-Parametric [453]. semiparametric [36]. semisupervised [38]. Senn [146]. Sensitivity [453, 112, 84]. separable [111]. sequence [315]. sequences [206]. Series [283, 590, 491, 167, 178, 110, 53, 121, 345, 75, 259]. Session [412, 405, 406, 425, 430, 417, 415, 411, 418, 404, 423, 402, 420, 416, 403, 427, 410, 429, 407, 428, 408, 409, 426, 419, 424]. Setting [579]. Settings [412, 401, 204, 18]. Severini [268]. severity [309]. Shafer [138, 128, 129, 127, 139, 132, 136, 140, 141, 142, 143, 144, 145, 133, 130, 134, 131, 146, 151, 137, 148, 135, 149, 150, 147]. Shahjahan [280]. Shao [368]. Shaoran [357, 361]. shares [511]. Sharing [442, 313]. SHELF [175]. Shi [351]. shiny [187]. shipping [57]. ships [117]. short [575]. Shrinkage [478]. Shuyi [357, 361]. side [5]. sided [28]. Signal [269]. signs [28, 30]. similarity [45]. similarity-based [45]. Simon [274]. Simple [44, 173]. simulated [241]. Singh [364]. Single [343, 173]. Single-month [343]. Sir [490, 393]. six [296]. size [547, 152, 233]. sizes [30]. skewed [304]. skills [114]. Skinner [92, 333]. slave [311]. slavery [263, 42]. Small [455, 450, 546, 545, 347, 178, 65, 12, 180, 14, 7]. Smartphone [435]. Smith [67, 333, 535]. soccer [577]. Social [552, 211, 440, 472, 580, 556, 260, 524, 13, 518, 243, 271, 200, 369, 519, 292, 303, 85, 568, 523]. society [193, 513, 196, 412, 405, 406, 425, 417, 415, 411, 418, 404, 423, 402, 420, 416, 403, 427, 410, 429, 407, 428, 408, 409, 426, 419, 424]. socio [32, 60, 383, 54]. socio-economic [32, 60, 383, 54]. Solari [536]. Solutions [587]. Solvency [588]. Sons [271, 284]. South [572, 314]. Sovereign [434]. Space [491, 525, 237, 343, 116, 523]. space-time [237]. Spain [256, 394]. Sparse [382]. Spatial [97, 553, 9, 55, 419, 413, 34, 227, 312, 275, 486, 549, 308, 516, 58, 567, 235, 182, 39, 374]. Spatial-Temporal [97]. spatially [545]. Spatio [197, 452]. Spatio-Temporal [452, 197]. spatiotemporal [313]. Speaking [325]. Special [412, 405, 406, 425, 430, 417, 415, 411, 418, 404, 423, 402, 420, 416, 403, 427, 410, 429, 407, 428, 408, 409, 426, 419, 424, 333]. specific [229]. Specification [105]. spillover [57]. spillovers [56]. spline [312]. Sports [268]. Spread [422, 243]. Springer [512, 285, 280]. SPSS [271]. stable [73]. stage [335]. stakes [208]. Stan [491, 245]. Standard [412, 401, 579, 29]. Starters [20]. Stata [272]. State [491, 340, 343, 116]. State-of-the [340]. State-Space [491, 343]. statements [305]. stationary [48]. Statistical [349, 326, 352, 353, 322, 493, 412, 354, 405, 99, 355, 406, 425, 356, 417, 357, 415, 167, 562, 358, 411, 418, 404, 423, 23, 359, 360, 218, 402, 361, 282, 362, 363, 420, 416, 403, 427, 410, 429, 421, 407, 428, 408, 409, 426, 351, 364, 348, 419, 424, 323, 350, 365, 30, 481, 138, 128, 129, 127, 247, 47, 139, 132, 240, 136, 140, 141, 142, 143, 144, 145, 59, 133, 130, 134, 131, 146, 151, 126, 137, 148, 135, 149, 150, 147, 368, 366, 367, 249, 98, 245, 274, 285, 559, 191]. statistician [591]. Statistics [587, 221, 330, 475, 586, 558, 268, 246, 529, 247, 271, 324, 290, 116, 279, 157, 332, 331, 489, 466, 468, 222, 122, 467]. Stensrud [500]. Stephen [146]. Steven [277, 409, 426]. stochastic [526, 232]. Stone [119]. stop [316]. strata [372, 334]. Strategies [492, 571]. strategy [138, 128, 129, 127, 139, 132, 136, 140, 141, 142, 143, 144, 145, 133, 130, 134, 131, 146, 151, 126, 137, 148, 135, 149, 311, 150, 147]. street [316]. street-level [316]. streets [195]. strength [50]. Strengths [161]. stress [73]. Stronger [78]. Structural [14, 110, 544, 54]. Structure [433]. Structure-Preserving [433]. structures [88]. student [114, 82, 71, 375]. student-teacher [71]. Studies [442, 453, 312, 172, 29, 16, 70, 73, 569, 280, 462]. Study [300, 295, 50, 205, 78, 161, 71, 87, 293, 295]. style [8]. subgroups [569]. subjective [163, 472]. subnational [374]. sucks [571]. summer [159]. superlative [158]. superspreading [381]. Supervised [469]. Supplement [300]. supply [107, 5]. supply-side [5]. support [255, 571, 515, 565]. surge [316]. surgical [81]. surveillance [547, 198]. Survey [443, 440, 444, 432, 439, 296, 441, 294, 301, 232, 371, 298, 299, 485, 176, 335, 115, 345, 302, 576, 343, 14]. Surveys [442, 34, 13, 299, 113, 292, 569]. Survival [474, 16, 104]. Sustainable [585]. swimmers [118]. Swiss [14]. Switching [300]. Switzerland [512, 285]. Sylvia [331, 408, 332]. syndrome [314]. synthesis [339]. Synthetic [113, 517, 183]. System [451, 531, 538, 533, 255, 536, 539, 534, 543, 530, 540, 537, 202, 541, 313, 515, 532, 535, 542, 233]. Systemic [451]. systems [263, 42].

T [275, 512, 278]. T. [67]. Tales [581]. Tao [350]. Targeting [512]. tax [152, 62]. Taylor [282, 277, 374]. teacher [71]. Techniques [581]. technology [574]. telephone [300]. Telescope [242]. temperature [48]. temperatures [234]. Temporal [521, 97, 62, 374, 452, 382, 550, 39, 197]. term [182]. terrorism [201]. terrorist [515]. Terry [445]. Test [269]. Test-Based [269]. Testing [138, 128, 129, 127, 139, 132, 238, 136, 140, 141, 142, 143, 144, 145, 133, 130, 134, 131, 146, 151, 126, 137, 148, 135, 149, 150, 147, 208, 573, 105]. tests [375, 317, 28]. Text [469, 198]. thanks [531, 331, 349, 128, 495, 127, 496, 532, 348]. Theo [593]. Theory [555, 492, 25, 118, 259, 284]. therapy [81]. thinking [326]. Thomas [268]. three [11, 195]. thresholds [173, 202]. Ties [244]. tilting [6]. Tim [487]. Time [590, 491, 429, 421, 345, 310, 237, 242, 178, 288, 48, 577, 110, 53, 16, 375, 75, 117, 380, 259, 347, 121]. Time-Series [491]. time-to-event [310, 380]. time-varying [242, 16]. times [257, 331, 330, 332]. Tiwari [269]. Tobias [166]. Toby [166]. Tomba [424]. Tony [489]. too [412, 401]. tool [258, 196]. Tools [454, 580, 279]. top [511]. Topic [412, 405, 406, 425, 430, 417, 415, 411, 418, 404, 423, 402, 420, 416, 403, 427, 410, 429, 407, 428, 408, 409, 426, 419, 424]. topological [517]. Trace [436, 442]. track [198]. Tracking [443, 8, 277]. trade [311]. Training [446, 228]. traits [514, 315]. transatlantic [311]. transfers [111]. transformations [7, 342]. transit [262]. Transition [300]. Transitioning [294]. transitions [54]. Transmission [412, 405, 406, 425, 430, 417, 415, 411, 418, 404, 423, 402, 420, 416, 403, 427, 410, 429, 407, 428, 408, 409, 426, 419, 424]. Transnational [306]. transport [43]. transportation [274]. trapdoor [180]. traumatic [73]. treated [112]. treatment [489, 307, 228, 112, 212, 570, 61, 154, 17, 15]. treatment-induced [15]. treatments [106, 242, 73, 219]. trends [178, 76, 229, 39, 198]. trial [37, 10]. trial-based [37]. Trials [99, 562, 1, 253, 84, 154, 380]. triangulation [517]. Trustworthiness [240]. truthfulness [305]. Twin [264]. twins [264]. Twitter [441]. Two [204, 335, 28, 181, 442]. two-level [335]. Two-phase [204]. two-sided [28]. two-stage [335]. Tyler [507]. types [52, 181]. Tze [353, 142].

Uganda [210, 56]. UK [511, 6, 62, 565].

uncertainty [331, 330, 117, 332].

uncounted [281]. Understand [268, 154]. Understanding [436, 271, 30, 489, 13]. unemployment [343]. Unified [188]. unifying [514]. Union [5]. unique [264]. unit [160, 377]. unit-level [377]. United [420, 255, 414, 207]. units [78]. Univariate [586, 271, 279]. unreliable [153]. Unveiling [548, 39]. unweighted [372]. Updating [433]. uplift [512]. Upper [108]. upside [57]. urban [11]. USA [271]. usage [314]. USD [273, 276]. Use [440, 301, 174, 411, 400]. user [262, 339]. user-friendly [339]. Using [372, 299, 444, 339, 435, 433, 439, 589, 559, 241, 71, 475, 586, 558, 268, 198, 207, 310, 526, 376, 32, 1, 253, 307, 572, 232, 344, 546, 98, 343, 295, 485, 60, 6, 484, 79, 14, 206, 346, 64, 75, 370, 379, 279, 118, 229, 569, 175, 523, 156, 233]. utilization [3]. Uttar [39].

validation [204, 370]. validity [66]. value [62, 81, 118]. value-added [62]. values [233]. VanderWeele’s [507]. variable [205, 79, 80, 83]. variables [169, 180, 346]. variance [179]. variation [549, 115]. various [52]. vary [48]. varying [242, 483, 36, 16]. varying-coefficient [36]. vector [6, 575]. Venice [548]. versus [76]. Via [437, 304, 90, 346, 8, 191]. Victor [505]. View [556]. vignettes [105]. violation [373]. violence [257, 373]. Violet [385]. Visualization [187]. Visualizations [123]. Vladimir [135]. Vlas [412]. volatility [544].

Volume [479, 102, 225, 396, 573]. Vos [148]. vote [531, 331, 349, 128, 495, 127, 496, 532, 348]. Vovk [135]. vulnerability [211].

W [269, 387, 123, 282]. wage [551]. Walter [215]. Wan [365]. Wang [149, 350]. warning [202]. Washington [274]. Wasserstein [550]. water [547, 91]. water-borne [547]. waves [296]. weak [50]. weapons [111]. weather [227]. Web [187, 294, 301, 300, 443, 298, 295, 437]. Web-Based [187]. web-first [300]. weeks [347]. Wei [366, 365]. weight [73]. Weighting [338, 112, 66]. weighting-based [112]. weights [1, 253, 157]. Well [472, 163]. Well-Being [472, 163]. Weng [354]. Wetherill [319]. Where [571]. Whittle [250]. Who [449, 489]. Wijayatunga [150, 508]. Wiley [271, 284, 279]. Willem [94]. William [214]. Wishart [216]. within [577, 5]. women [39]. Wong [354]. Woolford [542]. working [289, 290]. world [260, 463]. wound [81]. wounds [81]. Written [596, 297].

Xiao [133]. Xiao-Li [133]. Xiaoping [351]. xii [556, 281]. xiii [282]. xv [277]. xvii [285, 279]. xxi [271, 560]. xxx [583]. xxxiv [284].

Yingwei [283]. young [228]. Yu [283, 559]. Yucheng [351]. Yue [351].

Zalkikar [269]. Zealand [241, 233]. zero [174]. Zhang [531, 538, 533, 536, 539, 534, 543, 540, 537, 541, 351, 532, 535, 542]. Zhao [368, 367]. Zhong [365]. Zobitz [589]. Zou [365]. Zurich [422]. Zwet [94].

References

Buchanan:2018:GER

[1] Ashley L. Buchanan, Michael G. Hudgens, Stephen R. Cole, Katie R. Mollan, Paul E. Sax, Eric S. Daar, Adaora A. Adimora, Joseph J. Eron, and Michael J. Mugavero. Generalizing evidence from randomized trials using inverse probability of sampling weights. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4): 1193–1209, October 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See corrigendum [253].

Anonymous:2020:IIa

[2] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183 (1):1–2, January 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Yoon:2020:EGH

[3] Sungwook Yoon, Duk Bin Jun, and Sungho Park. The effect of general health checks on healthcare utilization: accounting for self-selection bias. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1): 3–36, January 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chambers:2020:ISA

[4] Ray Chambers and Andrea Diniz da Silva. Improved secondary analysis of linked data: a framework and an illustration. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1):37–59, January 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Meertens:2020:DDS

[5] Q. A. Meertens, C. G. H. Diks, H. J. van den Herik, and F. W. Takes. A data-driven supply-side approach for estimating cross-border Internet purchases within the European Union. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1): 61–90, January 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Koop:2020:URN

[6] Gary Koop, Stuart McIntyre, and James Mitchell. UK regional nowcasting using a mixed frequency vector autoregressive model with entropic tilting. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183 (1):91–119, January 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Rojas-Perilla:2020:DDT

[7] Natalia Rojas-Perilla, Sören Pannier, Timo Schmid, and Nikos Tzavidis. Data-driven transformations in small area estimation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1):121–148, January 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ross:2020:TEL

[8] Gordon J. Ross. Tracking the evolution of literary style via Dirichletmultinomial change point regression. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1):

149–167, January 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Paci:2020:SHM

[9] Lucia Paci, Alan E. Gelfand, María Asunción Beamonte, Pilar Gargallo, and Manuel Salvador. Spatial hedonic modelling adjusted for preferential sampling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1): 169–192, January 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Rhodes:2020:ATR

[10] K. M. Rhodes, J. Savović, R. Elbers, H. E. Jones, J. P. T. Higgins, J. A. C. Sterne, N. J. Welton, and R. M. Turner. Adjusting trial results for biases in metaanalysis: combining data-based evidence on bias with detailed trial assessment. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1): 193–209, January 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Arbabi:2020:PIU

[11] Hadi Arbabi, Martin Mayfield, and Philip McCann. Productivity, infrastructure and urban density — an allometric comparison of three European city regions across scales. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1): 211–228, January 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Frumento:2020:PMQ

[12] Paolo Frumento and Nicola Salvati. Parametric modelling of M-quantile regression coefficient functions with application to small area estimation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1): 229–250, January 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dawkins:2020:WDC

[13] Laura C. Dawkins, Daniel B. Williamson, Stewart W. Barr, and Sally R. Lampkin. ‘What drives commuter behaviour?’: a Bayesian clustering approach for understanding opposing behaviours in social surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1):251–280, January 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Molina:2020:EPS

[14] Isabel Molina and Ewa Strzalkowska-Kominiak. Estimation of proportions in small areas: application to the labour force using the Swiss Census Structural Survey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1):281–310, January 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Wodtke:2020:RRE

[15] Geoffrey T. Wodtke, Zahide Alaca, and Xiang Zhou. Regression-with-residuals estimation of marginal effects: a method of adjusting for treatment-induced confounders that may also be effect modifiers. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1):311–332, January 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lambert:2020:ITV

[16] Philippe Lambert and Vincent Bremhorst. Inclusion of time-varying covariates in cure survival models with an application in fertility studies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1):333–354, January 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Vegetabile:2020:OBG

[17] Brian G. Vegetabile, Daniel L. Gillen, and Hal S. Stern. Optimally balanced Gaussian process propensity scores for estimating treatment effects. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1): 355–377, January 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

McGee:2020:ODS

[18] Glen McGee, Jonathan Schildcrout, Sharon-Lise Normand, and Sebastien Haneuse. Outcome-dependent sampling in cluster-correlated data settings with application to hospital profiling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1): 379–402, January 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2020:BRa

[19] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183 (1):403, January 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kumar:2020:BRR

[20] Kuldeep Kumar. Book review: Regression Analysis in Medical Research for Starters and 2nd Levelers. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183 (1):404, January 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Papalia:2020:BRF

[21] Rosa Bernardini Papalia. Book review: Foundations of Info-metrics: Modeling, Inference, and Imperfect Information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1): 404–405, January 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

King:2020:BRP

[22] Thomas King. Book review: Practical R for Mass Communication and Journalism. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1):405, January 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kumar:2020:BRC

[23] Kuldeep Kumar. Book review: Classification, (Big) Data Analysis and Statistical Learning. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1):405–406, January 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Aalabaf-Sabaghi:2020:BRM

[24] Morteza Aalabaf-Sabaghi. Book review: A Multivariate Claim Count Model for Applications in Insurance.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 183 (1):406, January 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Reese:2020:BRP

[25] R. Allan Reese. Book review: Philogenetic Inference, Selection Theory, and History of Science. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1):406–407, January 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hopkins:2020:BRB

[26] Johnny Hopkins. Book review: The Beauty of Mathematics in Computer Science. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1):407, January 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2020:IIb

[27] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(2): 409–410, February 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Rice:2020:KSD

[28] Kenneth Rice, Tyler Bonnett, and Chloe Krakauer. Knowing the signs: a direct and generalizable motivation of twosided tests. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(2):411–430, February 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Held:2020:NSA

[29] Leonhard Held. A new standard for the analysis and design of replication studies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183 (2):431–448, February 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2020:DMS

[30] Anonymous. Discussion on the meeting on ‘Signs and sizes:understanding and replicating statistical findings’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(2): 449–469, February 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Heaps:2020:IEP

[31] Sarah E. Heaps, Malcolm Farrow, and Kevin J. Wilson. Identifying the effect of public holidays on daily demand for gas. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(2): 471–492, February 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bijlsma:2020:MSE

[32] Maarten J. Bijlsma and Ben Wilson. Modelling the socio-economic determinants of fertility: a mediation analysis using the parametric g-formula. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(2): 493–513, February 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Wadsworth:2020:QFI

[33] Ian Wadsworth, Lisa V. Hampson, Thomas Jaki, Graeme J. Sills, Anthony G. Marson, and Richard Appleton. A quantitative framework to inform extrapolation decisions in children. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(2): 515–534, February 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bakar:2020:BSC

[34] K. Shuvo Bakar, Nicholas Biddle, Philip Kokic, and Huidong Jin. A Bayesian spatial categorical model for prediction to overlapping geographical areas in sample surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(2):535–563, February 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Otting:2020:HHP

[35] Marius Ötting, Roland Langrock, Christian Deutscher, and Vianey Leos-Barajas. The hot hand in professional darts. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(2):565–580, February 2020. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Jin:2020:IAL

[36] Man Jin, Shunan Zhao, and Subal C. Kumbhakar. Information asymmetry and leverage adjustments: a semiparametric varying-coefficient approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(2): 581–605, February 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gabrio:2020:BPA

[37] Andrea Gabrio, Michael J. Daniels, and Gianluca Baio. A Bayesian parametric approach to handle missing longitudinal outcome data in trial-based health economic evaluations. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(2):607–629, February 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Li:2020:IOR

[38] Zhiyong Li, Xinyi Hu, Ke Li, Fanyin Zhou, and Feng Shen. Inferring the outcomes of rejected loans: an application of semisupervised clustering. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(2): 631–654, February 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Vicente:2020:CAW

[39] G. Vicente, T. Goicoa, P. Fernandez-Rasines, and M. D. Ugarte. Crime against women in India: unveiling spatial patterns and temporal trends of dowry deaths in the districts of Uttar Pradesh. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(2):655–679, February 2020. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Munafo:2020:CGR

[40] Marcus Munafò, Neil M. Davies, and George Davey Smith. Can genetics reveal the causes and consequences of educational attainment? Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(2): 681–688, February 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2020:IIc

[41] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3): 689–690, June 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Silverman:2020:MSA

[42] Bernard W. Silverman. Multiplesystems analysis for the quantification of modern slavery: classical and Bayesian approaches. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3):691–736, June 2020. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ouali:2020:GDP

[43] Laila Ait Bihi Ouali, Daniel J. Graham, Alexander Barron, and Mark Trompet. Gender differences in the perception of safety in public transport. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3): 737–769, June 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Jung:2020:SRG

[44] Jongbin Jung, Connor Concannon, Ravi Shroff, Sharad Goel, and Daniel G. Goldstein. Simple rules to guide expert classifications. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3):771–800, June 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dendramis:2020:SBA

[45] Y. Dendramis, G. Kapetanios, and M. Marcellino. A similarity-based approach for macroeconomic forecasting. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183 (3):801–827, June 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ellison:2020:FCF

[46] Joanne Ellison, Erengul Dodd, and Jonathan J. Forster. Forecasting of cohort fertility under a hierarchical Bayesian approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3):829–856, June 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Finazzi:2020:FIN

[47] Francesco Finazzi. Fulfilling the information need after an earthquake: statistical modelling of citizen science seismic reports for predicting earthquake parameters in near realtime. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3): 857–882, June 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dagsvik:2020:HDT

[48] John K. Dagsvik, Mariachiara Fortuna, and Sigmund Hov Moen. How does temperature vary over time?: evidence on the stationary and fractal nature of temperature fluctuations. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3): 883–908, June 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Fagan:2020:CPA

[49] Brennen T. Fagan, Marina I. Knight, Niall J. MacKay, and A. Jamie Wood. Change point analysis of historical battle deaths. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3):909–933, June 2020. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Heng:2020:FSW

[50] Siyu Heng, Dylan S. Small, and Paul R. Rosenbaum. Finding the strength in a weak instrument in a study of cognitive outcomes produced by Catholic high schools. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3):935–958, June 2020. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Elfadaly:2020:QEO

[51] Fadlalla G. Elfadaly and Paul H. Garthwaite. On quantifying expert opinion about multinomial models that contain covariates. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3):959–981, June 2020. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Brown:2020:NAL

[52] Sarah Brown, William Greene, and Mark Harris. A novel approach to latent class modelling: identifying the various types of body mass index individuals. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183 (3):983–1004, June 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Galbraith:2020:QAB

[53] Christopher Galbraith, Padhraic Smyth, and Hal S. Stern. Quantifying the association between discrete event time series with applications to digital forensics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3):1005–1027, June 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Zhu:2020:MSE

[54] Yajing Zhu, Fiona Steele, and Irini Moustaki. A multilevel structural equation model for the interrelationships between multiple latent dimensions of childhood socio-economic circumstances, partnership transitions and mid-life health. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3):1029–1050, June 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Pereira:2020:SCH

[55] João B. M. Pereira, Widemberg S. Nobre, Igor F. L. Silva, and Alexandra M. Schmidt. Spatial confounding in hurdle multilevel beta models: the case of the Brazilian Mathematical Olympics for Public Schools. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3): 1051–1073, June 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Wilke:2020:PDD

[56] Anna M. Wilke, Donald P. Green, and Jasper Cooper. A placebo design to detect spillovers from an educationentertainment experiment in Uganda.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3): 1075–1096, June 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Theodossiou:2020:FRD

[57] Panayiotis Theodossiou, Dimitris Tsouknidis, and Christos Savva. Freight rates in downside and upside markets: pricing of own and spillover risks from other shipping segments. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3):1097–1119, June 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Keller:2020:SSS

[58] Joshua P. Keller and Adam A. Szpiro. Selecting a scale for spatial confounding adjustment. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3):1121–1143, June 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Mathur:2020:NSM

[59] Maya B. Mathur and Tyler J. VanderWeele. New statistical metrics for multisite replication projects. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3): 1145–1166, June 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kjaergaard:2020:LFS

[60] Søren Kjærgaard, Yunus Emre Ergemen, Marie-Pier Bergeron-Boucher, Jim Oeppen, and Malene Kallestrup-Lamb. Longevity forecasting by socio-economic groups using compositional data analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183 (3):1167–1187, June 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Phillippo:2020:MNM

[61] David M. Phillippo, Sofia Dias, A. E. Ades, Mark Belger, Alan Brnabic, Alexander Schacht, Daniel Saure, Zbigniew Kadziola, and Nicky J. Welton. Multilevel network meta-regression for population-adjusted treatment comparisons. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3):1189–1210, June 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Labonne:2020:TDO

[62] Paul Labonne and Martin Weale. Temporal disaggregation of overlapping noisy quarterly data: estimation of monthly output from UK value-added tax data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3):1211–1230, June 2020. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Scott:2020:MBC

[63] Marc A. Scott, Kaushik Mohan, and Jacques-Antoine Gauthier. Model-based clustering and analysis of life history data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3):1231–1251, June 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Nethery:2020:CIF

[64] Rachel C. Nethery, Yue Yang, Anna J. Brown, and Francesca Dominici. A causal inference framework for cancer cluster investigations using publicly available data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3):1253–1272, June 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Fabrizi:2020:FAS

[65] Enrico Fabrizi, Maria Rosaria Ferrante, and Carlo Trivisano. A functional approach to small area estimation of the relative median poverty gap. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3): 1273–1291, June 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Wang:2020:IEV

[66] Lingxiao Wang, Barry I. Graubard, Hormuzd A. Katki, and Yan Li. Improving external validity of epidemiologic cohort analyses: a kernel weighting approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3):1293–1311, June 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2020:MFS

[67] Anonymous. T. M. F. Smith, 1934–2019; Harvey Goldstein, 1939–2020; Allan Henry Seheult, 1942–2019; John Francis Bithell, 1939–2020; M. H. A. Davis, 1945–2020. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3):1313–1322, June 2020. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2020:BRb

[68] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(3): 1323–1326, June 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2020:IId

[69] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4): 1327–1328, October 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Mattei:2020:PPC

[70] Alessandra Mattei, Bianca L. De Stavola, and Fabrizia Mealli. Preface to the papers on ‘Causal inference from non-experimental studies: challenges, developments and applications’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4): 1329–1332, October 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lockwood:2020:UHI

[71] J. R. Lockwood and D. McCaffrey. Using hidden information and performance level boundaries to study studentteacher assignments: implications for estimating teacher causal effects. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4): 1333–1362, October 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Longford:2020:PAA

[72] Nicholas T. Longford. Performance assessment as an application of causal inference. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1363–1385, October 2020. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Resa:2020:DSW

[73] María de los Angeles Resa and José R. Zubizarreta. Direct and stable weight adjustment in non-experimental studies with multivalued treatments: analysis of the effect of an earthquake on posttraumatic stress. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1387–1410, October 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Pimentel:2020:OMA

[74] Samuel D. Pimentel, Lauren Vollmer Forrow, Jonathan Gellar, and Jiaqi Li. Optimal matching approaches in health policy evaluations under rolling enrolment. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1411–1435, October 2020. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Samartsidis:2020:BMF

[75] Pantelis Samartsidis, Shaun R. Seaman, Silvia Montagna, André Charlett, Matthew Hickman, and Daniela De Angelis. A Bayesian multivariate factor analysis model for evaluating an intervention by using observational time series data on multiple outcomes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4): 1437–1459, October 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hunt:2020:BVG

[76] Lamar Hunt III, Irene B. Murimi, Jodi B. Segal, Marissa J. Seamans, Daniel O. Scharfstein, and Ravi Varadhan. Brand versus generic: addressing non-adherence, secular trends and nonoverlap. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1461–1478, October 2020. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

vonCube:2020:CIM

[77] Maja von Cube, Martin Schumacher, and Martin Wolkewitz. Causal inference with multistate models-estimands and estimators of the population attributable fraction. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1479–1500, October 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Keele:2020:SIR

[78] Luke Keele, Steve Harris, Samuel D. Pimentel, and Richard Grieve. Stronger instruments and refined covariate balance in an observational study of the effectiveness of prompt admission to intensive care units. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1501–1521, October 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Mauro:2020:IVM

[79] Jacqueline A. Mauro, Edward H. Kennedy, and Daniel Nagin. Instrumental variable methods using dynamic interventions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1523–1551, October 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Sanchez:2020:FIV

[80] Guillermo Briseño Sanchez, Maike Hohberg, Andreas Groll, and Thomas Kneib. Flexible instrumental variable distributional regression. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4): 1553–1574, October 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Saramago:2020:BEM

[81] Pedro Saramago, Karl Claxton, Nicky J. Welton, and Marta Soares. Bayesian econometric modelling of observational data for cost-effectiveness analysis: establishing the value of negative pressure wound therapy in the healing of open surgical wounds. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1575–1593, October 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Licari:2020:ACE

[82] Federica Licari and Alessandra Mattei. Assessing causal effects of extra compulsory learning on college students’ academic performances. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4): 1595–1614, October 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Zeng:2020:BOC

[83] Shuxi Zeng, Fan Li, and Peng Ding. Is being an only child harmful to psychological health?: evidence from an instrumental variable analysis of China’s one-child policy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1615–1635, October 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Imai:2020:ISA

[84] Kosuke Imai and Zhichao Jiang. Identification and sensitivity analysis of contagion effects in randomized placebocontrolled trials. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1637–1657, October 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ogburn:2020:CIS

[85] Elizabeth L. Ogburn, Ilya Shpitser, and Youjin Lee. Causal inference, social networks and chain graphs. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4): 1659–1676, October 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kim:2020:HEP

[86] Chanmin Kim, Lucas R. F. Henneman, Christine Choirat, and Corwin M. Zigler. Health effects of power plant emissions through ambient air quality. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4): 1677–1703, October 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Andrews:2020:ECM

[87] Ryan M. Andrews, Ilya Shpitser, Oscar Lopez, William T. Longstreth, Paulo H. M. Chaves, Lewis Kuller, and Michelle C. Carlson. Examining the causal mediating role of brain pathology on the relationship between diabetes and cognitive impairment: the Cardiovascular Health Study. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1705–1726, October 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Castelletti:2020:DCS

[88] Federico Castelletti and Guido Consonni. Discovering causal structures in Bayesian Gaussian directed acyclic graph models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1727–1745, October 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Foraita:2020:CDG

[89] Ronja Foraita, Juliane Friemel, Kathrin Günther, Thomas Behrens, Jörn Bullerdiek, Rolf Nimzyk, Wolfgang Ahrens, and Vanessa Didelez. Causal discovery of gene regulation with incomplete data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1747–1775, October 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Corradi:2020:CEB

[90] Fabio Corradi and Monica Musio. Causes of effects via a Bayesian model selection procedure. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1777– 1792, October 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Cuellar:2020:NPP

[91] Maria Cuellar and Edward H. Kennedy. A non-parametric projection-based estimator for the probability of causation, with application to water sanitation in Kenya. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1793–1818, October 2020. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chambers:2020:CJS

[92] Ray Chambers, Ian Diamond, Tim Holt, Jouni Kuha, Danny Pfeffermann, Natalie Shlomo, Pedro Nascimento Silva, Paul Smith, David Steel, and Fiona Steele. Christopher John Skinner, 1953– 2020. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1819–1821, October 2020. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chatfield:2020:GJG

[93] Chris Chatfield. Gerald Joseph Goodhardt, 1930–2020. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1821–1823, October 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bickel:2020:WVZ

[94] P. J. Bickel and N. I. Fisher. Willem van Zwet, 1934–2020. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1823–1825, October 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2020:BR

[95] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4): 1827, October 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Aalabaf-Sabaghi:2020:BRI

[96] Morteza Aalabaf-Sabaghi. Book review: An Introduction to Data, Everything You Need to Know About AI: Big Data and Data Science. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1828, October 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chaturvedi:2020:BRM

[97] Anoop Chaturvedi. Book review: Modelling Spatial and Spatial-Temporal Data: a Bayesian Approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4): 1828–1829, October 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dietz:2020:BRS

[98] Sebastian Dietz. Book review: Statistical Data Analysis using SAS. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4): 1829, October 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

DiTanna:2020:BRM

[99] Gian Luca Di Tanna. Book review: Modern Adaptive Randomized Clinical Trials: Statistical and Practical Aspects. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4): 1829–1830, October 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2020:R

[100] Anonymous. Referees. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1831– 1836, October 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2020:IA

[101] Anonymous. Index of authors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4): 1837–1844, October 2020. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2020:CV

[102] Anonymous. Contents of volume 183, 2020. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(4):1845–1848, October 2020. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2021:IIe

[103] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184 (1):1–2, January 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Tanner:2021:DSP

[104] Kamaryn T. Tanner, Linda D. Sharples, Rhian M. Daniel, and Ruth H. Keogh. Dynamic survival prediction combining landmarking with a machine learning ensemble: Methodology and empirical comparison. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1):3–30, January 2021. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Greene:2021:STH

[105] William H. Greene, Mark N. Harris, Rachel J. Knott, and Nigel Rice. Specification and testing of hierarchical ordered response models with anchoring vignettes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1):31–64, January 2021. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bansak:2021:ECM

[106] Kirk Bansak. Estimating causal moderation effects with randomized treatments and non-randomized moderators. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184 (1):65–86, January 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Cai:2021:EHE

[107] Lixin Cai. The effects of health on the extensive and intensive margins of labour supply. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1):87–117, January 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Christensen:2021:MIC

[108] Bent Jesper Christensen, Nabanita Datta Gupta, and Paolo Santucci de Magistris. Measuring the impact of clean energy production on CO_{2} abatement in Denmark: Upper bound estimation and forecasting. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1):118–149, January 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Crossley:2021:IEM

[109] Thomas F. Crossley, Tobias Schmidt, Panagiota Tzamourani, and Joachim K. Winter. Interviewer effects and the measurement of financial literacy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1): 150–178, January 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dorsett:2021:BST

[110] Richard Dorsett. A Bayesian structural time series analysis of the effect of basic income on crime: Evidence from the Alaska Permanent Fund. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1): 179–200, January 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lebacher:2021:DSN

[111] Michael Lebacher, Paul W. Thurner, and Göran Kauermann. A dynamic separable network model with actor heterogeneity: an application to global weapons transfers. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1):201–226, January 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hong:2021:DYC

[112] Guanglei Hong, Fan Yang, and Xu Qin.

Did you conduct a sensitivity analysis? A new weighting-based approach for evaluations of the average treatment effect for the treated. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1):227–254, January 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kim:2021:SME

[113] Hang J. Kim, Jörg Drechsler, and Katherine J. Thompson. Synthetic microdata for establishment surveys under informative sampling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1): 255–281, January 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Knaus:2021:DML

[114] Michael C. Knaus. A double machine learning approach to estimate the effects of musical practice on student’s skills. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1): 282–300, January 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Moore:2021:DCV

[115] Jamie C. Moore, Gabriele B. Durrant, and Peter W. F. Smith. Do coefficients of variation of response propensities approximate non-response biases during survey data collection? Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1): 301–323, January 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Schiavoni:2021:DFM

[116] Caterina Schiavoni, Franz Palm, Stephan Smeekes, and Jan van den Brakel. A dynamic factor model approach to incorporate Big Data in state space models for official statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1): 324–353, January 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Song:2021:PDT

[117] Yunting Song and Nuo Wang. On probability distributions of the time deviation law of container liner ships under interference uncertainty. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1):354–367, January 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Spearing:2021:ROP

[118] Harry Spearing, Jonathan Tawn, David Irons, Tim Paulden, and Grace Bennett. Ranking, and other properties, of elite swimmers using extreme value theory. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1): 368–395, January 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Galbraith:2021:MS

[119] Rex Galbraith. Mervyn Stone, 1932– 2020. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1):396–398, January 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Goodall:2021:FJ

[120] Gerald Goodall. Flavia Jolliffe 1942– 2020. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1):398–399, January 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kumar:2021:BRT

[121] Kuldeep Kumar. Book review: Time series clustering and classification. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184 (1):400, January 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kumar:2021:BRS

[122] Kuldeep Kumar. Book review: Statistics and Health Care Fraud: How to Save Billions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1):401, January 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dietz:2021:BRR

[123] Sebastian Dietz. Book review: R Visualizations — Derive Meaning from Data, Gerbing, D. W., 2020. Chapman and Hall/CRC Press, Boca Raton. 237 pp. 68,99 GBP. ISBN 978-1-138-59963-5. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184 (1):401–402, January 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dietz:2021:BRH

[124] Sebastian Dietz. Book review: Handbook of Financial Risk Management. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1): 402–403, January 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2021:IIf

[125] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2): 405–406, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shafer:2021:TBS

[126] Glenn Shafer. Testing by betting: a strategy for statistical and scientific communication. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):407–431, April 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See discussion [138, 128, 129, 127, 139, 132, 136, 140, 141, 142, 143, 144, 145, 133, 130, 134, 131, 146, 137, 148, 135, 149, 150, 147] and reply [151].

Dawid:2021:PVT

[127] Philip Dawid. Proposer of the vote of thanks to Glenn Shafer and contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):432–433, April 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Coolen:2021:SVT

[128] Frank P. A. Coolen. Seconder of the vote of thanks to Glenn Shafer and contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184 (2):434–435, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Crane:2021:HCC

[129] Harry Crane. Harry Crane’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 184 (2):436–437, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Osimani:2021:BOC

[130] Barbara Osimani. Barbara Osimani’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):437–438, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Ramdas:2021:ARC

[131] Aaditya Ramdas. Aaditya Ramdas’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):438–440, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Grunwald:2021:PDG

[132] Peter D. Grünwald. Peter D. Grünwald’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):440–441, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Meng:2021:XLM

[133] Xiao-Li Meng. Xiao-Li Meng’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184 (2):442–443, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Pedersen:2021:APP

[134] Arthur Paul Pedersen. Arthur Paul Pedersen’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):443–444, April 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Vovk:2021:VVC

[135] Vladimir Vovk. Vladimir Vovk’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):445–446, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Hennig:2021:CHC

[136] Christian Hennig. Christian Hennig’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):446–447, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Smith:2021:DLC

[137] Paul A. Smith. A distillation of the live chat during the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):448, April 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Chai:2021:CPC

[138] Christine P. Chai. Christine P. Chai’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):449–450, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Greenland:2021:SGC

[139] Sander Greenland. Sander Greenland’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):450–451, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Krakauer:2021:CKK

[140] Chloe Krakauer and Kenneth Rice. Chloe Krakauer and Kenneth Rice’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):452–453, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Kumar:2021:KKC

[141] Kuldeep Kumar. Kuldeep Kumar’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):453–454, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Lai:2021:TLL

[142] Tze Leung Lai and Anna Choi. Tze Leung Lai and Anna Choi’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184 (2):454–455, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Longford:2021:NLC

[143] Nick Longford. Nick Longford’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):455–456, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Martin:2021:RMC

[144] Ryan Martin. Ryan Martin’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 184 (2):456–457, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Mateu:2021:JMC

[145] Jorge Mateu. Jorge Mateu’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):458, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Senn:2021:SSC

[146] Stephen Senn. Stephen Senn’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184 (2):459–460, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

terSchure:2021:JTS

[147] Judith ter Schure. Judith ter Schure’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):460–461, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Vos:2021:PVC

[148] Paul Vos. Paul Vos’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2): 461–462, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Wang:2021:RWC

[149] Ruodu Wang. Ruodu Wang’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184 (2):463–464, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Wijayatunga:2021:PWC

[150] Priyantha Wijayatunga. Priyantha Wijayatunga’s contribution to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2): 465–466, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Shafer:2021:ARD

[151] Glenn Shafer. Author’s reply to the Discussion of ‘Testing by betting: a strategy for statistical and scientific communication’ by Glenn Shafer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2): 466–478, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [126].

Dellaportas:2021:SSD

[152] Petros Dellaportas, Evangelos Ioannidis, and Christos Kotsogiannis. Sample size determination for risk-based tax auditing. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):479–493, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Rougier:2021:EER

[153] Jonathan Rougier. Estimating eventrates from unreliable historical records. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184 (2):494–503, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Tipton:2021:BGA

[154] Elizabeth Tipton. Beyond generalization of the ATE: Designing randomized trials to understand treatment effect heterogeneity. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):504–521, April 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Zhang:2021:LDL

[155] Li-Chun Zhang and Tiziana Tuoto. Linkage-data linear regression. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2): 522–547, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

vanBerkum:2021:QLG

[156] Frank van Berkum, Katrien Antonio, and Michel Vellekoop. Quantifying longevity gaps using micro-level lifetime data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):548–570, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Zhang:2021:PEW

[157] Li-Chun Zhang. Proxy expenditure weights for Consumer Price Index: Audit sampling inference for big-data statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):571–588, April 2021. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

vonAuer:2021:CAS

[158] Ludwig von Auer and Jochen Wengenroth. Consistent aggregation with superlative and other price indices. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2): 589–615, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

McNeish:2021:SDM

[159] Daniel McNeish and Denis Dumas. A seasonal dynamic measurement model for summer learning loss. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2): 616–642, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Akande:2021:LAI

[160] Olanrewaju Akande, Gabriel Madson, D. Sunshine Hillygus, and Jerome P. Reiter. Leveraging auxiliary information on marginal distributions in nonignorable models for item and unit nonresponse. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):643–662, April 2021. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Livingston:2021:ENF

[161] Mark Livingston, Francesca Pannullo, Adrian W. Bowman, E. Marian Scott, and Nick Bailey. Exploiting new forms of data to study the private rented sector: Strengths and limitations of a database of rental listings. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):663–682, April 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Proietti:2021:NMG

[162] Tommaso Proietti and Alessandro Giovannelli. Nowcasting monthly GDP with big data: a model averaging approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184 (2):683–706, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kavetsos:2021:EBR

[163] Georgios Kavetsos, Ichiro Kawachi, Ilias Kyriopoulos, and Sotiris Vandoros. The effect of the Brexit referendum result on subjective well-being. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2): 707–731, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Weidmann:2021:MPD

[164] Ben Weidmann and Luke Miratrix. Missing, presumed different: Quantifying the risk of attrition bias in education evaluations. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2):732–760, April 2021. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Doucouliagos:2021:HAG

[165] Chris Doucouliagos, Jack Hennessy, and Debdulal Mallick. Health aid, governance and infant mortality. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2): 761–783, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Barnett:2021:TTL

[166] Vic Barnett and Kevin McConway. Tobias (Toby) Lewis 1918–2020. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(2): 784–786, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gomez-Rubio:2021:BRH

[167] Virgilio Gómez-Rubio. Book review: Handbook of Mixture Analysis, Frühwirth Schnatter, S., Celeux, G. and Robert, C.P., 2019. Chapman and Hall/ CRC Handbooks of Modern Statistical Methods Series, Boca Raton. 522 pp., 52.99 GBP. ISBN 978 0 367 732066. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184 (2):787–788, April 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2021:IIg

[168] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3):

789–790, July 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Aliverti:2021:RIG

[169] Emanuele Aliverti, Kristian Lum, James E. Johndrow, and David B. Dunson. Removing the influence of group variables in high-dimensional predictive modelling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3):791–811, July 2021. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Fe:2021:PER

[170] Eduardo Fé. Pension eligibility rules and the local causal effect of retirement on cognitive functioning. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3): 812–841, July 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Fannon:2021:MNL

[171] Zoë Fannon, Christiaan Monden, and Bent Nielsen. Modelling non-linear age-period-cohort effects and covariates, with an application to English obesity 2001–2014. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3):842–867, July 2021. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hedges:2021:DRS

[172] Larry V. Hedges and Jacob M. Schauer. The design of replication studies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3): 868–886, July 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Katki:2021:SFI

[173] Hormuzd A. Katki and Ionut Bebu. A simple framework to identify optimal cost-effective risk thresholds for a single screen: Comparison to Decision Curve Analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3):887–903, July 2021. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Erosheva:2021:WZM

[174] Elena A. Erosheva, Patrícia Martinková, and Carole J. Lee. When zero may not be zero: a cautionary note on the use of inter-rater reliability in evaluating grant peer review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3):904–919, July 2021. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Williams:2021:CPE

[175] Cameron J. Williams, Kevin J. Wilson, and Nina Wilson. A comparison of prior elicitation aggregation using the classical method and SHELF. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184 (3):920–940, July 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kim:2021:CNP

[176] Jae Kwang Kim, Seho Park, Yilin Chen, and Changbao Wu. Combining non-probability and probability survey samples through mass imputation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184 (3):941–963, July 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Jorda:2021:IMG

[177] Vanesa Jorda, José María Sarabia, and Markus Jäntti. Inequality measurement with grouped data: Parametric and non-parametric methods. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3): 964–984, July 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Boonstra:2021:MTS

[178] Harm Jan Boonstra, Jan van den Brakel, and Sumonkanti Das. Multilevel time series modelling of mobility trends in the Netherlands for small domains. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3): 985–1007, July 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dong:2021:LVE

[179] Ziqing Dong, Yves Tillé, Giovanni M. Giorgi, and Alessio Guandalini. Linearization and variance estimation of the Bonferroni inequality index. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3): 1008–1029, July 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Helske:2021:ECE

[180] Jouni Helske, Santtu Tikka, and Juha Karvanen. Estimation of causal effects with small data in the presence of trapdoor variables. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3):1030–1051, July 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Silber:2021:EQR

[181] Henning Silber, Joss Roßmann, Tobias Gummer, Stefan Zins, and Kai Willem Weyandt. The effects of question, respondent and interviewer characteristics on two types of item nonresponse. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3): 1052–1069, July 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Schliep:2021:LTS

[182] Erin M. Schliep, Alan E. Gelfand, Jesús Abaurrea, Jesús Asín, María A. Beamonte, and Ana C. Cebrián. Longterm spatial modelling for characteristics of extreme heat events. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3): 1070–1092, July 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Quick:2021:GPD

[183] Harrison Quick. Generating Poissondistributed differentially private synthetic data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3):1093–1108, July 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kunz:2021:PIE

[184] Johannes S. Kunz, Kevin E. Staub, and Rainer Winkelmann. Predicting individual effects in fixed effects panel probit models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3):1109–1145, July 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Goldie:2021:JH

[185] Charles M. Goldie. John Haigh 1941– 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3):1146–1148, July 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

DeStavola:2021:MHJ

[186] Bianca L. De Stavola. Michael Hills (7 June 1934-7 January 2021). Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3):1149, July 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2021:BRI

[187] Shalabh. Book review: Interactive Web-Based Data Visualization with R, plotly, and shiny. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3):1150, July 2021. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

French:2021:BRE

[188] Simon French. Book review: The Equation of Knowledge: From Bayes’ Rule to a Unified Philosophy of Science. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3): 1151–1152, July 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Jolliffe:2021:BRD

[189] Ian Jolliffe. Book review: Dark data: Why what you don’t know matters.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3):1153, July 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chaturvedi:2021:BRA

[190] Anoop Chaturvedi. Book review: Analyzing high-dimensional gene expression and DNA methylation data with R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3):1154, July 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2021:BRS

[191] Shalabh. Book review: Statistical inference via data science. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3):1155, July 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2021:IIh

[192] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4): 1157–1158, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hardelid:2021:DSS

[193] Pia Hardelid, Peter Christen, Elizabeth Williamson, Katie Harron, and Bianca L. De Stavola. Data science for society: Challenges, developments and applications. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1159–1160, October 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Meng:2021:EPD

[194] Xiao-Li Meng. Enhancing (publications on) data quality: Deeper data minding and fuller data confession. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4): 1161–1175, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bolin:2021:FAM

[195] David Bolin, Vilhelm Verendel, Meta Berghauser Pont, Ioanna Stavroulaki, Oscar Ivarsson, and Erik Håkansson. Functional ANOVA modelling of pedestrian counts on streets in three European cities. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4): 1176–1198, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Wright:2021:PNC

[196] Marvin N. Wright, Sasmita Kusumastuti, Laust H. Mortensen, Rudi G. J. Westendorp, and Thomas A. Gerds. Personalised need of care in an ageing society: the making of a prediction tool based on register data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4): 1199–1219, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Virtanen:2021:STM

[197] Seppo Virtanen and Mark Girolami. Spatio-temporal mixed membership models for criminal activity. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4): 1220–1244, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

You:2021:UTM

[198] Jingxian You, Paul Expert, and Céire Costelloe. Using text mining to track outbreak trends in global surveillance of emerging diseases: ProMED-mail. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4): 1245–1259, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Geroldinger:2021:EPC

[199] Angelika Geroldinger, Milan Hronsky, Florian Endel, Gottfried Endel, Rainer Oberbauer, and Georg Heinze. Estimation of the prevalence of chronic kidney disease in people with diabetes by combining information from multiple routine data collections. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1260–1282, October 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Iacopini:2021:FIP

[200] Matteo Iacopini and Carlo R. M. A.

Santagiustina. Filtering the intensity of public concern from social media count data with jumps. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1283–1302, October 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Tickle:2021:CEH

[201] S. O. Tickle, I. A. Eckley, and P. Fearnhead. A computationally efficient, highdimensional multiple changepoint procedure with application to global terrorism incidence. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1303–1325, October 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Masselot:2021:MLA

[202] Pierre Masselot, Fateh Chebana, Céline Campagna, Éric Lavigne, Taha B. M. J. Ouarda, and Pierre Gosselin. Machine learning approaches to identify thresholds in a heat-health warning system context. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1326–1346, October 2021. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Fu:2021:MPR

[203] Qiang Fu, Tian-Yi Zhou, and Xin Guo. Modified Poisson regression analysis of grouped and right-censored counts. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4): 1347–1367, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Amorim:2021:TPS

[204] Gustavo Amorim, Ran Tao, Sarah Lotspeich, Pamela A. Shaw, Thomas Lumley, and Bryan E. Shepherd. Twophase sampling designs for data validation in settings with covariate measurement error and continuous outcome. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4): 1368–1389, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hocagil:2021:PSA

[205] Tugba Akkaya Hocagil, Richard J. Cook, Sandra W. Jacobson, Joseph L. Jacobson, and Louise M. Ryan. Propensity score analysis for a semi-continuous exposure variable: a study of gestational alcohol exposure and childhood cognition. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1390–1413, October 2021. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Murphy:2021:CLL

[206] Keefe Murphy, T. Brendan Murphy, Raffaella Piccarreta, and I. Claire Gormley. Clustering longitudinal life-course sequences using mixtures of exponentialdistance models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1414–1451, October 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

vanDijk:2021:ULC

[207] Justin T. van Dijk, Guy Lansley, and Paul A. Longley. Using linked consumer registers to estimate residential moves in the United Kingdom. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4): 1452–1474, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Adhikari:2021:ALA

[208] Samrachana Adhikari, Tracy Sweet, and Brian Junker. Analysis of longitudinal advice-seeking networks following implementation of high stakes testing. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4): 1475–1500, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Walter:2021:DPG

[209] Paul Walter, Marcus Groß, Timo Schmid, and Nikos Tzavidis. Domain prediction with grouped income data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4): 1501–1523, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Egami:2021:CSG

[210] Naoki Egami and Erin Hartman. Covariate selection for generalizing experimental results: Application to a largescale development program in Uganda. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4): 1524–1548, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Didkovskyi:2021:SMV

[211] Oleksandr Didkovskyi, Giovanni Azzone, Alessandra Menafoglio, and Piercesare Secchi. Social and material vulnerability in the face of seismic hazard: an analysis of the Italian case. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4): 1549–1577, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lee:2021:PPP

[212] Youjin Lee, Trang Q. Nguyen, and Elizabeth A. Stuart. Partially pooled propensity score models for average treatment effect estimation with multilevel data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4): 1578–1598, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Page:2021:RDE

[213] Eric Page and David A. Elston. Roger David Elston (1928–2021). Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1599– 1600, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Mallon:2021:WGS

[214] John Mallon, Vivienne Stevenson, and Gilbert MacKenzie. William Gilmore Stevenson (1943–2021). Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1600– 1601, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Horn:2021:KKR

[215] Stephen Horn, Raymond Chambers, and William Gross. Kenneth (Ken) Ronald Walter Brewer (1931–2021). Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1602– 1604, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Barker:2021:DW

[216] Sarah Barker and Shirley Coleman.

David Wishart (1943–2020). Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1604– 1605, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Trewin:2021:LL

[217] Dennis Trewin. Lars Lyberg, (1944– 2021). Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1606–1607, October 2021. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lazic:2021:BRC

[218] Stanley E. Lazic. Book review: A Computational Approach to Statistical Learning. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1608, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Jen:2021:BRB

[219] Min-Hua Jen. Book review: Bayesian cost-effectiveness analysis of medical treatments. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1608–1609, October 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dietz:2021:BRP

[220] Sebastian Dietz. Book review: Probability Companion for Engineering and Computer Science. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1609–1610, October 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gepp:2021:BRA

[221] Adrian Gepp. Book review: Advanced Statistics with Applications in R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4): 1610–1611, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hossain:2021:BRS

[222] Md Moyazzem Hossain. Book review: Statistics for Data Science and Policy Analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1612, October 2021. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2021:BRR

[223] Shalabh. Book review: R Markdown Cookbook. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1613, October 2021. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2021:R

[224] Anonymous. Referees. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1614– 1621, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2021:CV

[225] Anonymous. Contents of volume 184, 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1622–1625, October 2021. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2022:IIa

[226] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (1):1–2, January 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bille:2022:EWC

[227] Anna Gloria Billé and Marco Rogna.

The effect of weather conditions on fertilizer applications: a spatial dynamic panel data analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1):3–36, January 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See correction [328].

Dorsett:2022:PAT

[228] Richard Dorsett and Lucy Stokes. Preapprenticeship training for young people: Estimating the marginal and average treatment effects. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1): 37–60, January 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Stefanucci:2022:ACS

[229] Marco Stefanucci and Stefano Mazzuco. Analysing cause-specific mortality trends using compositional functional data analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1):61–83, January 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Deb:2022:EME

[230] Soudeep Deb and Sougata Deb. An ensemble method for early prediction of dengue outbreak. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1):84–101, January 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Karmakar:2022:IRA

[231] Bikram Karmakar, Peng Liu, Gourab Mukherjee, Hai Che, and Shantanu Dutta. Improved retention analysis in freemium role-playing games by jointly modelling players’ motivation, progression and churn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1):102–133, January 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Cernat:2022:ESS

[232] Alexandru Cernat and Daniel L. Oberski. Estimating stochastic survey response errors using the multitraitmultierror model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1):134–155, January 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

vanderHeijden:2022:MSE

[233] Peter G. M. van der Heijden, Maarten Cruyff, Paul A. Smith, Christine Bycroft, Patrick Graham, and Nathaniel Matheson-Dunning. Multiple system estimation using covariates having missing values and measurement error: Estimating the size of the Maori population in New Zealand. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1):156–177, January 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bhargava:2022:EMC

[234] Alok Bhargava. Econometric modelling of carbon dioxide emissions and concentrations, ambient temperatures and ocean deoxygenation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1): 178–201, January 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Python:2022:DAC

[235] Andre Python, Andreas Bender, Marta Blangiardo, Janine B. Illian, Ying Lin, Baoli Liu, Tim C. D. Lucas, Siwei Tan, Yingying Wen, Davit Svanidze, and Jianwei Yin. A downscaling approach to compare COVID-19 count data from databases aggregated at different spatial scales. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1):202–218, January 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lombardia:2022:NAG

[236] María José Lombardía, Esther López-Vizcaíno, and Cristina Rueda. A new approach to the gender pay gap decomposition by economic activity. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1): 219–245, January 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bartolucci:2022:HMS

[237] Francesco Bartolucci and Alessio Farcomeni. A hidden Markov spacetime model for mapping the dynamics of global access to food. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1): 246–266, January 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hannig:2022:TCD

[238] Jan Hannig and Hari Iyer. Testing for calibration discrepancy of reported likelihood ratios in forensic science. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1): 267–301, January 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shi:2022:AHR

[239] Peng Shi, Glenn M. Fung, and Daniel Dickinson. Assessing hail risk for property insurers with a dependent marked point process. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1):302–328, January 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hand:2022:TSI

[240] David J. Hand. Trustworthiness of statistical inference. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1):329–347, January 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lee:2022:UMS

[241] Sanghyeok Lee and Tue Gørgens. Using maximum simulated likelihood methods to overcome left censoring: Dynamic event history models of heart attack risk in New Zealand. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1):348–376, January 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Blackwell:2022:TMR

[242] Matthew Blackwell and Anton Strezhnev. Telescope matching for reducing model dependence in the estimation of the effects of time-varying treatments: an application to negative advertising. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1): 377–399, January 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Fritz:2022:IRM

[243] Cornelius Fritz and Göran Kauermann. On the interplay of regional mobility, social connectedness and the spread of COVID-19 in Germany. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1): 400–424, January 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lazic:2022:BRM

[244] Stanley E. Lazic. Book review: Medical Risk Prediction Models: With Ties to Machine Learning. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1):425, January 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Green:2022:BRS

[245] Nathan Green. Book review: Statistical Rethinking: A Bayesian Course with Examples in R and Stan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1): 426–427, January 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Banks:2022:BRR

[246] David Banks. Book review: Review of statistics and the evaluation of evidence for forensic scientists. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1): 428–429, January 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dienes:2022:BRE

[247] Zoltan Dienes. Book review: Evidencebased statistics: an introduction to the evidential approach-from likelihood principle to statistical practice. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (1):430, January 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kumar:2022:BRM

[248] Kuldeep Kumar. Book review: Measuring abundance. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1):431, January 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Caporin:2022:BRS

[249] Massimiliano Caporin. Book review:

Statistical Analysis of Financial Data: with Examples In R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1):432– 433, January 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kelly:2022:PW

[250] Frank Kelly. Peter Whittle, 1927–2021.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1): 434–437, January 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Davis:2022:RPJ

[251] Tim Davis. Richard Parry-Jones 1951– 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1):438–440, January 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Capell:2022:JMM

[252] Peter Capell. John Michael (Mike) Haslam 1947–2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1):441–442, January 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Buchanan:2022:CGE

[253] Ashley L. Buchanan and Michael G. Hudgens. Corrigendum: Generalizing evidence from randomized trials using inverse probability of sampling weights. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1): 443–444, January 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [1].

Anonymous:2022:IIb

[254] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2): 445–446, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Barons:2022:DSS

[255] Martine J. Barons, Thais C. O. Fonseca, Andy Davis, and Jim Q. Smith. A decision support system for addressing food security in the United Kingdom. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):447–470, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Pavia:2022:ECEa

[256] Jose M. Pavía and Josep Lledó. Estimation of the combined effects of ageing and seasonality on mortality risk: an application to Spain. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):471–497, April 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anderberg:2022:QDV

[257] Dan Anderberg, Helmut Rainer, and Fabian Siuda. Quantifying domestic violence in times of crisis: an Internet search activity-based measure for the COVID-19 pandemic. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2): 498–518, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Floridi:2022:PCS

[258] Ginevra Floridi and Benjamin E. Lauderdale. Pairwise comparisons as a scale development tool for composite measures. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):519–542, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Zhang:2022:ATP

[259] Xinyu Zhang and Howell Tong. Asymptotic theory of principal component analysis for time series data with cautionary comments. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):543–565, April 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Clark:2022:CRW

[260] Duncan A. Clark and Mark S. Handcock. Comparing the real-world performance of exponential-family random graph models and latent order logistic models for social network analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (2):566–587, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lum:2022:CTT

[261] Kristian Lum, David B. Dunson, and James Johndrow. Closer than they appear: a Bayesian perspective on individual-level heterogeneity in risk assessment. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):588–614, April 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bansal:2022:DCM

[262] Prateek Bansal, Daniel Hörcher, and Daniel J. Graham. A dynamic choice model to estimate the user cost of crowding with large-scale transit data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2): 615–639, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Binette:2022:RMS

[263] Olivier Binette and Rebecca C. Steorts. On the reliability of multiple systems estimation for the quantification of modern slavery. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):640–676, April 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

vandenBerg:2022:UBT

[264] Gerard J. van den Berg and Bettina Drepper. A unique bond: Twin bereavement and lifespan associations of identical and fraternal twins. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2): 677–698, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ahundjanov:2022:PLC

[265] Behzod B. Ahundjanov, Sherzod B.

Akhundjanov, and Botir B. Okhunjanov. Power law in COVID-19 cases in China. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):699–719, April 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Morgan:2022:PMN

[266] Byron J. T. Morgan. Philip M. North 1949–2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):720–721, April 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Goldie:2022:DRR

[267] Charles M. Goldie. Derek Richard Robinson 1947–2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):722–723, April 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Yip:2022:BRA

[268] Stan Yip. Book review: Analytic Methods in Sports — Using Mathematics and Statistics to Understand Data from Baseball, Football, Basketball, and Other, Thomas A. Severini, 2020 CRC Press, ix + 372 pp., ISBN 978-0-367-46938-2. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):724, April 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chaturvedi:2022:BRS

[269] Anoop Chaturvedi. Book review: Signal Detection for Medical Scientists: Likelihood Ratio Based Test-Based Methodology, Ram C. Tiwari, Jyoti Zalkikar and Lan Zalkikar, 2021 Boca Roton, FL, Chapman & Hall/CRC Press 244 pp, 14 B/W Illustrations £89.99 (Hardbound), £40.49 (e-book) ISBN 978-0-367-20143-2. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (2):725–726, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Eyles:2022:BRB

[270] Emily Eyles. Book review: Beyond Multiple Linear Regression: Applied Generalized Linear Models and Multilevel Models in R, Paul Roback and Julie Legler 2021. 1st edn. Boca Raton, CRC Press. 418 pp. £74.95. ISBN: 978-1-4398-8538-3. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):726–727, April 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hossain:2022:BRA

[271] Md. Moyazzem Hossain. Book review: Applied univariate, bivariate, and multivariate statistics: Understanding statistics for social and natural scientists, with applications in SPSS and R, Daniel J. Denis, 2021 Hoboken, NJ, USA,, John Wiley & Sons, Inc. xxi + 551 pp., $104.00 (e-book), $130.00 (hardcover), ISBN: 978-1-119-58304-2. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2): 727–728, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chowdhry:2022:BRA

[272] Amit K. Chowdhry. Book review of Applied Meta-Analysis with R and Stata, Ding-Geng (Din) Chen and Karl E. Peace, 2021 Pages: 456, Price: $120, ISBN 13: 978-0-367-18383-7. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2): 728–729, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chen:2022:BRB

[273] Li-Pang Chen. Book review: Bayesian analysis of infectious diseases: COVID-19 and beyond, Lyle D. Broemeling, 2021 331 pages, USD $111.85 (hardback), ISBN 0-367-63386-8. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2): 729–730, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hewson:2022:BRS

[274] Paul Hewson. Book review: Statistical and econometric methods for transportation data analysis, Simon Washington, Matthew Karlaftis, Fred Mannering, Panagiotis Anastasopoulis 2020 Chapman Hall/CRC Press, 496 pp., ISBN 978-0-367-19902-9. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):731, April 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chaturvedi:2022:BRA

[275] Anoop Chaturvedi. Book review: Analyzing spatial models of choice and judgment, David A. Armstrong II, Ryan Bakker, Royce Carroll, Christopher Hare, Keith T. Poole, Howard Rosenthal, Chapman and Hall/CRC 2020 320 Pages, Hardback: £59.99, eBook: £53.99, ISBN 978-1-138-71533-2. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2): 732–733, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chen:2022:BRIa

[276] Li-Pang Chen. Book review: Introduction to data science: Data analysis and prediction algorithms with R, Chapman and Hall/CRC 2019, 743 pages. USD $92.63 (hardback). ISBN 978-0-367-35798-6. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):733–734, April 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Saunders:2022:BRM

[277] Catherine L. Saunders. Book review: Monitoring the health of populations by tracking disease outbreaks: Saving Humanity from the Next Plague, Steven E. Rigdon and Ronald D. Fricker, Jr., Boca Raton, FL, CRC Press, Taylor and Frances Group, xv + 200 pp. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2): 734–735, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2022:BRM

[278] Shalabh. Book review: Modern data science with R, Banjamin S. Baumer, Daniel T. Kaplan, Nicholas J. Horten, Chapman and Hall/CRC Press 214 ISBN 978-0-367-19149-8. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):735–736, April 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2022:BRU

[279] Shalabh. Book review: Univariate, bivariate and multivariate statistics using R: Quantitative tools for data analysis and data science, Daniel J. Denis, 2020 Hobokon, NJ, Wiley, ISBN 978-1-119-54993-2, xvii + 366. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):736–737, April 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Saleh:2022:BRM

[280] A. K. Md. Ehsanes Saleh. Book review:

Meta-analysis methods for health and experimental studies, Shahjahan Khan, 2020, 1st edition Springer Nature, CRC Press, 293 pages, Price 67,59 EUR (hardcopy) 53,49 EUR (eBook), ISBN 978-981-15-5032-4. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):737–738, April 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

King:2022:BRU

[281] Thomas King. Book review: The uncounted, Alex Cobham, Polity Press, 2019 xii + 227 pp. ISBN: 978-1-5095-3602-3, £14.99 (pbk) £45 (hbk). Journal of the Royal Statistical Society.

Series A (Statistics in Society), 185 (2):739, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Little:2022:RCA

[282] Mark P. Little. Review of A Computational Approach to Statistical Learning, Taylor Arnold, Michael Kane and Bryan W. Lewis, 2019, Chapman & Hall/CRC Boca Raton, FL xiii + 361, $110.00 ISBN 13 978-1-138-04637-5. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2): 740–741, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bullement:2022:BRC

[283] Ash Bullement. Book review: Cure models: Methods, applications and implementation, Yingwei Peng, Binbing Yu, 2021, Hardback, first edition, Chapman & Hall/CRC Biostatistics Series, 268 pages, £74.99, ISBN 978-0-367-14557-6. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (2):741–742, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2022:BRT

[284] Shalabh. Book review: Theory of ridge regression estimation with applications, A. K. Md. Ehsanes Saleh, Mohammad Arashi, B. M. Golam Kibria, 2019 John Wiley & Sons, Inc., Hoboken, pp. xxxiv + 342, ISBN 978-1-118-64461-4. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):742–743, April 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hossain:2022:BRS

[285] Md. Moyazzem Hossain. Book review:

Statistical regression modeling with R, Ding-Geng (Din) Chen, Jenny K. Chen, 2021, Springer Nature Switzerland AG xvii + 228 pp., £71.50 (eBook), £89.99 (hardcover) ISBN 9788-3-030-67582-0, ISBN 978-3-030-67583-7 (eBook). Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2): 743–744, April 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2022:IIc

[286] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3): 745–746, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Carpenter:2022:CHG

[287] James R. Carpenter. A celebration of Harvey Goldstein’s lifetime contributions: Foreword. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):747–748, July 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Cole:2022:CHG

[288] T. J. Cole. A celebration of Harvey Goldstein’s lifetime contributions: Harvey Goldstein and his time at the Institute of Child Health. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3): 748–752, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Browne:2022:CHG

[289] William J. Browne. A celebration of Harvey Goldstein’s lifetime contributions: Memories of working with Harvey Goldstein on multilevel modelling methods and applications. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3): 753–758, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Leckie:2022:CHG

[290] George Leckie. A celebration of Harvey Goldstein’s lifetime contributions: Memories of working with Harvey Goldstein on educational research and statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):758–762, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Harron:2022:CHG

[291] Katie Harron. A celebration of Harvey Goldstein’s lifetime contributions: a journey in data linkage with Harvey Goldstein. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):763–767, July 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Maslovskaya:2022:FOD

[292] Olga Maslovskaya, Bella Struminskaya, and Gabriele Durrant. The future of online data collection in social surveys: Challenges, developments and applications. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):768–772, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Cornesse:2022:GIP

[293] Carina Cornesse, Ulrich Krieger, Marie-Lou Sohnius, Marina Fikel, Sabine Friedel, Tobias Rettig, Alexander Wenz, Sebastian Juhl, Roni Lehrer, Katja Möhring, Elias Naumann, Maximiliane Reifenscheid, and Annelies G. Blom. From German Internet Panel to Mannheim Corona Study: Adaptable probability-based online panel infrastructures during the pandemic. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3): 773–797, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Biemer:2022:TPS

[294] Paul P. Biemer, Kathleen Mullan Harris, Brian J. Burke, Dan Liao, and Carolyn Tucker Halpern. Transitioning a panel survey from in-person to predominantly web data collection: Results and lessons learned. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):798–821, July 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Goodman:2022:IUW

[295] Alissa Goodman, Matt Brown, Richard J.

Silverwood, Joseph W. Sakshaug, Lisa Calderwood, Joel Williams, and George B. Ploubidis. The impact of using the Web in a mixed-mode follow-up of a longitudinal birth cohort study: Evidence from the National Child Development Study. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):822–850, July 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Maslovskaya:2022:RSW

[296] Olga Maslovskaya and Peter Lugtig. Representativeness in six waves of CROss-National Online Survey (CRONOS) panel. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):851–871, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gavras:2022:ICO

[297] Konstantin Gavras, Jan Karem Höhne, Annelies G. Blom, and Harald Schoen. Innovating the collection of open-ended answers: the linguistic and content characteristics of written and oral answers to political attitude questions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (3):872–890, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See correction [596].

Gleiser:2022:IWM

[298] Patrick Gleiser, Joseph W. Sakshaug, Marieke Volkert, Peter Ellguth, Susanne Kohaut, and Iris Möller. Introducing Web in a mixed-mode establishment survey: Effects on nonresponse. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (3):891–915, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gummer:2022:URS

[299] Tobias Gummer, Pablo Christmann, Sascha Verhoeven, and Christof Wolf. Using a responsive survey design to innovate self-administered mixed-mode surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):916–932, July 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Sastry:2022:STW

[300] Narayan Sastry and Katherine A. McGonagle. Switching from telephone to web-first mixed-mode data collection: Results from the Transition into Adulthood Supplement to the US Panel Study of Income Dynamics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3): 933–954, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bosch:2022:NEU

[301] Oriol J. Bosch, Melanie Revilla, Danish Daniel Qureshi, and Jan Karem Höhne. A new experiment on the use of images to answer web survey questions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):955–980, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Schwarz:2022:DPS

[302] Hannah Schwarz, Melanie Revilla, and Bella Struminskaya. Do previous survey experience and participating due to an incentive affect response quality? Evidence from the CRONOS panel. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (3):981–1003, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Menardi:2022:DBC

[303] Giovanna Menardi and Domenico De Stefano. Density-based clustering of social networks. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):1004–1029, July 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Aliverti:2022:DMM

[304] Emanuele Aliverti, Stefano Mazzuco, and Bruno Scarpa. Dynamic modelling of mortality via mixtures of skewed distribution functions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):1030–1048, July 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Yu:2022:MPC

[305] Qiushi Yu and Kevin M. Quinn. A multidimensional pairwise comparison model for heterogeneous perceptions with an application to modelling the perceived truthfulness of public statements on COVID-19. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):1049–1073, July 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Huber:2022:TML

[306] Martin Huber, David Imhof, and Rieko Ishii. Transnational machine learning with screens for flagging bid-rigging cartels. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):1074–1114, July 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Caron:2022:EIT

[307] Alberto Caron, Gianluca Baio, and Ioanna Manolopoulou. Estimating individual treatment effects using nonparametric regression models: a review.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (3):1115–1149, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gilardi:2022:MHA

[308] Andrea Gilardi, Jorge Mateu, Riccardo Borgoni, and Robin Lovelace. Multivariate hierarchical analysis of car crashes data considering a spatial network lattice. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):1150–1177, July 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Millimet:2022:CSN

[309] Daniel L. Millimet and Christopher F.

Parmeter. COVID-19 severity: a new approach to quantifying global cases and deaths. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):1178–1215, July 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Adeleke:2022:RDD

[310] Mariam O. Adeleke, Gianluca Baio, and Aidan G. O’Keeffe. Regression discontinuity designs for time-to-event outcomes: an approach using accelerated failure time models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):1216–1246, July 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Wiens:2022:MSE

[311] Ashton Wiens, Henry B. Lovejoy, Zachary Mullen, and Eric A. Vance. A modelling strategy to estimate conditional probabilities of African origins: the collapse of the Oyo Empire and the transatlantic slave trade, 1817–1836. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (3):1247–1270, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bobb:2022:ASC

[312] Jennifer F. Bobb, Maricela F. Cruz, Stephen J. Mooney, Adam Drewnowski, David Arterburn, and Andrea J. Cook. Accounting for spatial confounding in epidemiological studies with individuallevel exposures: an exposure-penalized spline approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):1271–1293, July 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Piter:2022:HBS

[313] Andreas Piter, Philipp Otto, and Hamza Alkhatib. The Helsinki bike-sharing system — insights gained from a spatiotemporal functional model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3): 1294–1318, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Park:2022:EEO

[314] Jinhwan Park, Duk Bin Jun, and Sungho Park. The effect of epidemic outbreak on healthcare usage: Lessons from the 2015 Middle East respiratory syndrome outbreak in South Korea. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3): 1319–1343, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Mencarini:2022:LCP

[315] Letizia Mencarini, Raffaella Piccarreta, and Marco Le Moglie. Life-course perspective on personality traits and fertility with sequence analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3): 1344–1369, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Braakmann:2022:DSS

[316] Nils Braakmann. Does stop and search reduce crime? Evidence from streetlevel data and a surge in operations following a high-profile crime. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3): 1370–1397, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Nikoloulopoulos:2022:OFC

[317] Aristidis K. Nikoloulopoulos. An onefactor copula mixed model for joint meta-analysis of multiple diagnostic tests. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):1398–1423, July 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ullah:2022:NMR

[318] Aman Ullah, Tao Wang, and Weixin Yao. Nonlinear modal regression for dependent data with application for predicting COVID-19. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):1424–1453, July 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Coleman:2022:GBW

[319] Shirley Coleman. G. Barrie Wetherill (1932–2022). Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):1454–1455, July 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Andersen:2022:NK

[320] Per Kragh Andersen. Niels Keiding (1944–2022). Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):1456–1457, July 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2022:NHG

[321] Anonymous. Nigel Hawkes: Gentleman health editor par excellence. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3): 1457–1458, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chowdhry:2022:SID

[322] Amit K. Chowdhry. Statistical issues in drug development. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):1459, July 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ugarte:2022:SLB

[323] María Dolores Ugarte. Statistical learning for big dependent data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):1460, July 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hossain:2022:PSN

[324] Md. Moyazzem Hossain. Practical statistics for nursing and health care. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):1461, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Holmes:2022:SAN

[325] Dawn Holmes. Speaking against number: Heidegger, language and the politics of calculation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):1462, July 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Cahusac:2022:STS

[326] Peter Cahusac. Statistical thinking from scratch: a primer for scientists. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):1463, July 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

French:2022:EMA

[327] Simon French. Explanatory model analysis: Explore, explain, and examine predictive models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(3):1464, July 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bille:2022:CBR

[328] Anna Gloria Billé and Marco Rogna.

Correction of Billé and Rogna (2021). Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (3):1465–1468, July 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [227].

Anonymous:2022:IId

[329] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4): 1469–1470, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Richardson:2022:STI

[330] Sylvia Richardson. Statistics in times of increasing uncertainty. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1471– 1496, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ashby:2022:PVT

[331] Deborah Ashby. Proposal of the vote of thanks for ‘Statistics in times of increasing uncertainty’, Sylvia Richardson’s Presidential Address. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1497– 1498, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Spiegelhalter:2022:DPA

[332] David Spiegelhalter. Discussion of Presidential address: Statistics in times of increasing uncertainty by Sylvia Richardson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1499–1500, October 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Smith:2022:PSI

[333] Paul A. Smith and Peter W. F. Smith. Preface to the Special Issue in memory of Fred Smith and Chris Skinner.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4): 1501–1503, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Fuller:2022:PSB

[334] Wayne A. Fuller. Post-strata based on sample quantiles. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1504–1521, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kim:2022:ACS

[335] Jae Kwang Kim, J. N. K. Rao, and Yonghyun Kwon. Analysis of clustered survey data based on two-stage informative sampling and associated two-level models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1522–1540, October 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Zhang:2022:SBD

[336] Li-Chun Zhang and Gustav Haraldsen.

Secure big data collection and processing: Framework, means and opportunities. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1541–1559, October 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

vandenBrakel:2022:EML

[337] Jan van den Brakel, Martijn Souren, and Sabine Krieg. Estimating monthly labour force figures during the COVID-19 pandemic in the Netherlands. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):

1560–1583, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chambers:2022:WIC

[338] Ray Chambers, Setareh Ranjbar, Nicola Salvati, and Barbara Pacini. Weighting, informativeness and causal inference, with an application to rainfall enhancement. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1584–1612, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Jackson:2022:USC

[339] James Jackson, Robin Mitra, Brian Francis, and Iain Dove. Using saturated count models for user-friendly synthesis of large confidential administrative databases. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1613–1643, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shlomo:2022:MRR

[340] Natalie Shlomo and Chris Skinner. Measuring risk of re-identification in microdata: State-of-the art and new directions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1644–1662, October 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Silva:2022:FMM

[341] Pedro Luis do N. Silva and Fernando Antônio da S. Moura. Fitting multivariate multilevel models under informative sampling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1663– 1678, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Wurz:2022:ERI

[342] Nora Würz, Timo Schmid, and Nikos Tzavidis. Estimating regional income indicators under transformations and access to limited population auxiliary information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1679–1706, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Goncalves:2022:SMU

[343] Caio Gonçalves, Luna Hidalgo, Denise Silva, and Jan van den Brakel. Singlemonth unemployment rate estimates for the Brazilian Labour Force Survey using state-space models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1707–1732, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Clark:2022:SDA

[344] Robert G. Clark and David G. Steel.

Sample design for analysis using highinfluence probability sampling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4): 1733–1756, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Pfeffermann:2022:TSM

[345] Danny Pfeffermann. Time series modelling of repeated survey data for estimation of finite population parameters. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):

1757–1777, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Nason:2022:QER

[346] Guy P. Nason and James L. Wei. Quantifying the economic response to COVID-19 mitigations and death rates via forecasting purchasing managers’ indices using generalised network autoregressive models with exogenous variables. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1778–1792, October 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Zhao:2022:SDB

[347] Qingyuan Zhao. Small data, big time — a retrospect of the first weeks of COVID-19. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (4):1793–1814, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Studley:2022:PVT

[348] Ruth Studley. Proposer of the vote of thanks and contribution to the Discussion of ‘Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1815–1816, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bird:2022:SVT

[349] Sheila M. Bird. Seconder of the vote of thanks and contribution to the ‘First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):

1817–1819, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Wang:2022:TWC

[350] Tao Wang. Tao Wang’s contribution to the ‘First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4): 1819–1821, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shi:2022:XSY

[351] Xiaoping Shi, Yue Zhang, and Yucheng Dong. Xiaoping Shi, Yue Zhang and Yucheng Dong’s contribution to the ‘First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4): 1821–1822, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chai:2022:CPC

[352] Christine P. Chai. Christine P.

Chai’s contribution to the ‘First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1823– 1824, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Choi:2022:ACT

[353] Anna L. Choi. Anna L. Choi and Tze Leung Lai’s contribution to the ‘First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4): 1824–1825, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Cook:2022:ARC

[354] Alex R. Cook, Kwok Pui Choi, and Weng Kee Wong. Alex R. Cook, Kwok Pui Choi and Weng Kee Wong’s contribution to the ‘First Discussion Meeting on Statistical Aspects of the COVID-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1826–1827, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Diggle:2022:PJDa

[355] Peter J. Diggle. Peter J. Diggle’s contribution to the ‘First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1827–1828, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Farcomeni:2022:AFM

[356] Alessio Farcomeni and Marco Geraci. Alessio Farcomeni and Marco Geraci’s contribution to the ‘First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1829– 1830, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ge:2022:SGO

[357] Shuyi Ge, Oliver Linton, and Shaoran Li. Shuyi Ge, Oliver Linton and Shaoran Li’s contribution to the ‘First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4): 1831–1832, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Jana:2022:KJL

[358] Kingsuk Jana, Lagnajita Basu, and Kaushik Jana. Kingsuk Jana, Lagnajita Basu and Kaushik Jana’s contribution to the ‘First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1832–1833, October 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kumar:2022:KKF

[359] Kuldeep Kumar. Kuldeep Kumar’s first contribution to the ‘First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1833– 1834, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kumar:2022:KKS

[360] Kuldeep Kumar. Kuldeep Kumar’s second contribution to the ‘First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1834– 1835, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Li:2022:SLO

[361] Shaoran Li, Oliver Linton, and Shuyi Ge. Shaoran Li, Oliver Linton and Shuyi Ge’s contribution to the ‘First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4): 1836–1837, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Mateu:2022:JMF

[362] Jorge Mateu. Jorge Mateu’s first contribution to the ‘First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1838–1839, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Mateu:2022:JMS

[363] Jorge Mateu. Jorge Mateu’s second contribution to the ‘First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1839–1840, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Singh:2022:GNS

[364] Garib Nath Singh. Garib Nath Singh’s contribution to the ‘First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1840– 1841, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Zhong:2022:WZC

[365] Wei Zhong, Chuang Wan, and Changliang Zou. Wei Zhong, Chuang Wan and Changliang Zou’s contribution to the ‘First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4): 1841–1843, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Nason:2022:NWR

[366] Guy P. Nason and James L. Wei. Nason and Wei’s reply to the Discussion of ‘The First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4): 1844–1846, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Zhao:2022:ZRD

[367] Qingyuan Zhao. Zhao’s reply to the Discussion of ‘The First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1846–1848, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Jiang:2022:JZS

[368] Feiyu Jiang, Zifeng Zhao, and Xiaofeng Shao. Jiang, Zhao and Shao’s reply to the Discussion of ‘The First Discussion Meeting on Statistical Aspects of the Covid-19 Pandemic’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4): 1849–1854, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Koskinen:2022:BAS

[369] Johan Koskinen and Galina Daraganova. Bayesian analysis of social influence. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4): 1855–1881, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Schenkel:2022:AMU

[370] Jonas F. Schenkel and Li-Chun Zhang. Adjusting misclassification using a second classifier with an external validation sample. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1882–1902, October 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gao:2022:NNR

[371] Chenyin Gao, Katherine Jenny Thompson, Jae Kwang Kim, and Shu Yang. Nearest neighbour ratio imputation with incomplete multinomial outcome in survey sampling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1903–1930, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Brumberg:2022:URR

[372] Katherine Brumberg, Dylan S. Small, and Paul R. Rosenbaum. Using randomized rounding of linear programs to obtain unweighted natural strata that balance many covariates. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4): 1931–1951, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Roychowdhury:2022:DCL

[373] Punarjit Roychowdhury and Gaurav Dhamija. Don’t cross the line: Bounding the causal effect of hypergamy violation on domestic violence in India. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4): 1952–1978, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Yang:2022:TST

[374] Yang Yang, Han Lin Shang, and Joel E.

Cohen. Temporal and spatial Taylor’s law: Application to Japanese subnational mortality rates. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):1979– 2006, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

McNeish:2022:MTM

[375] Daniel McNeish, Denis Dumas, Dario Torre, and Neil Rice. Modelling time to maximum competency in medical student progress tests. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):2007–2034, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Betz:2022:CLE

[376] Jennifer Betz, Maximilian Nagl, and Daniel Rösch. Credit line exposure at default modelling using Bayesian mixed effect quantile regression. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):

2035–2072, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gardini:2022:PIM

[377] Aldo Gardini, Enrico Fabrizi, and Carlo Trivisano. Poverty and inequality mapping based on a unit-level log-normal mixture model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):2073–2096, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Mulick:2022:BHM

[378] Amy R. Mulick, Shefali Oza, David Prieto-Merino, Francisco Villavicencio, Simon Cousens, and Jamie Perin. A Bayesian hierarchical model with integrated covariate selection and misclassification matrices to estimate neonatal and child causes of death. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4): 2097–2120, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Seufert:2022:MEA

[379] Jacqueline D. Seufert, Andre Python, Christoph Weisser, Elías Cisneros, Krisztina Kis-Katos, and Thomas Kneib. Mapping ex ante risks of COVID-19 in Indonesia using a Bayesian geostatistical model on airport network data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4): 2121–2155, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Williams:2022:OPP

[380] Nicholas Williams, Michael Rosenblum, and Iván Díaz. Optimising precision and power by machine learning in randomised trials with ordinal and timeto-event outcomes with an application to COVID-19. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):2156–2178, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Meagher:2022:AEC

[381] Joe Meagher and Nial Friel. Assessing epidemic curves for evidence of superspreading. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):2179–2202, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Mosley:2022:STD

[382] Luke Mosley, Idris A. Eckley, and Alex Gibberd. Sparse temporal disaggregation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):2203–2233, October 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Verkroost:2022:BMH

[383] Florianne C. J. Verkroost. A Bayesian multivariate hierarchical growth curve model to examine cumulative socioeconomic (dis)advantage among childless adults and parents. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):2234– 2276, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2022:RML

[384] Anonymous. Robert Michael Loynes (1935–2021). Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):2277–2278, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2022:VRC

[385] Anonymous. Violet Rosina Cane, 1916– 2008. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):2278–2281, October 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Trewin:2022:BM

[386] Dennis Trewin. Bill McLennan (1942– 2022). Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):2282–2284, October 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Cormack:2022:WFK

[387] Richard Cormack. A W (Freda) Kemp (1930–2022). Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):2284–2285, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Molenberghs:2022:PD

[388] Geert Molenberghs. Pierre Dagnelie (1933–2022). Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):2285–2286, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Eilers:2022:IDC

[389] Paul Eilers and Maria Durbàn. Iain David Currie (1946–2022). Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):2287– 2288, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Haws:2022:BAJ

[390] Tony Haws. Barry Anthony John Quirke (1941–2022). Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):2288–2289, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lauritzen:2022:OEB

[391] Steffen Lauritzen and Michael Sørensen.

Ole Eiler Barndorff-Nielsen, 1935–2022. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4): 2289–2291, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lin:2022:NRD

[392] Dennis K. J. Lin. Norman Richard Draper (1931–2022). Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):2292– 2294, October 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Davison:2022:SDC

[393] Anthony C. Davison, Valerie S. Isham, and Nancy M. Reid. Sir David Cox: 1924–2022. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):2295–2306, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Pavia:2022:ECEb

[394] Jose M. Pavía and Josep Lledó. Estimation of the combined effects of ageing and seasonality on mortality risk: an application to Spain. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):2307–2309, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Pavia:2022:R

[395] Jose M. Pavía and Josep Lledó. Referees. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (4):2310–2319, October 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2022:CVa

[396] Anonymous. Contents of volume 185, 2022. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(4):2320–2325, October 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2022:IIe

[397] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S1–S2, November 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        185/Supplement_1/S1/7069491.

Diggle:2022:I

[398] Peter J. Diggle and Sylvia Richardson. ‘Introduction’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S1):S3– S4, November 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/article/
185/Supplement_1/S3/7069498.

Parag:2022:EGR

[399] Kris V. Parag, Robin N. Thompson, and Christl A. Donnelly. Are epidemic growth rates more informative than reproduction numbers? Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S5–S15, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S5/7069470.

Jewell:2022:URN

[400] Nicholas P. Jewell and Joseph A. Lewnard. On the use of the reproduction number for SARS-COV-2: Estimation, misinterpretations and relationships with other ecological measures. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S16–S27, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S16/7069477.

Coffeng:2022:PEI

[401] Luc E. Coffeng and Sake J. Vlas. Predicting epidemics and the impact of interventions in heterogeneous settings: Standard SEIR models are too pessimistic. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S1):S28–S35, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S28/7069476.

Lessler:2022:JLC

[402] Justin Lessler and C. Jessica E. Metcalf. Justin Lessler and C. Jessica E. Metcalf’s invited discussion: Contribution to the papers in session 1 of the Royal Statistical Society’s special topic meeting on COVID-19 transmission: 9 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S1):S36–S38, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S36/7069497.

ONeill:2022:PDO

[403] Philip D. O’Neill. Philip D. O’Neill’s invited discussion: Contribution to the papers in session 1 of the Royal Statistical Society’s special topic meeting on Covid-19 transmission: 9 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S39–S40, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S39/7069473.

Kingman:2022:JKI

[404] John Kingman. John Kingman’s invited discussion: Contribution to the papers in session 1 of the Royal Statistical Society’s special topic meeting on COVID-19 transmission: 9 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S41–S43, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S41/7069482.

Dagpunar:2022:JDD

[405] John Dagpunar. John Dagpunar’s discussion contribution to papers in session 1 of the Royal Statistical Society’s special topic meeting on COVID-19 transmission: 9 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S44–S46, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S44/7069496.

Diggle:2022:PJDb

[406] Peter J. Diggle. Peter J. Diggle’s discussion: Contribution to papers in session 1 of the Royal Statistical Society’s special topic meeting on COVID-19 transmission: 9 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S47–S48, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S47/7069495.

Pellis:2022:LPD

[407] Lorenzo Pellis. Lorenzo Pellis’ discussion: Contribution to papers in session 1 of the Royal Statistical Society’s special topic meeting on COVID-19 transmission: 9 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S49–S50, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S49/7069472.

Richardson:2022:SRD

[408] Sylvia Richardson. Sylvia Richardson’s discussion: Contribution to papers in session 1 of the Royal Statistical Society’s special topic meeting on COVID-19 transmission: 9 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S51–S52, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S51/7069489.

Riley:2022:SRDa

[409] Steven Riley. Steven Riley’s discussion: Contribution to papers in session 1 of the Royal Statistical Society’s special topic meeting on COVID-19 transmission: 9 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S53–S54, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S53/7069494.

Parag:2022:ARD

[410] Kris V. Parag, Robin N. Thompson, and Christl A. Donnelly. Authors’ reply to the discussion of ‘Are Epidemic Growth Rates More Informative than Reproduction Numbers?’ by Parag et al. in session 1 of the Royal Statistical Society’s special topic meeting on COVID-19 transmission: 9 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S55–S60, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S55/7069493.

Jewell:2022:ARD

[411] Nicholas P. Jewell and Joseph A. Lewnard. Authors’ reply to the discussion of ‘On the use of the Reproduction Number for SARS-CoV-2: Estimation, Misinterpretations, and Relationships with other Ecological Measures’ by Jewell & Lewnard in session 1 of the Royal Statistical Society’s special topic meeting on COVID-19 transmission: 9 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S1):S61–S62, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S61/7069488.

Coffeng:2022:ARD

[412] Luc E. Coffeng and Sake J. Vlas. Authors’ reply to the discussion of ‘Predicting Epidemics and the Impact of Interventions in Heterogeneous Settings: Standard SEIR Models are too Pessimistic’ by Coffeng & de Vlas in session 1 of the Royal Statistical Society’s special topic meeting on COVID-19 transmission: 9 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S63–S64, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S63/7069487.

Teh:2022:EBI

[413] Yee Whye Teh, Bryn Elesedy, Bobby He, Michael Hutchinson, Sheheryar Zaidi, Avishkar Bhoopchand, Ulrich Paquet, Nenad Tomasev, Jonathan Read, and Peter J. Diggle. Efficient Bayesian inference of instantaneous reproduction numbers at fine spatial scales, with an application to mapping and nowcasting the Covid-19 epidemic in British local authorities. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S1):S65–S85, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S65/7069481.

Mishra:2022:CML

[414] Swapnil Mishra, James A. Scott, Daniel J. Laydon, Harrison Zhu, Neil M. Ferguson, Samir Bhatt, Seth Flaxman, and Axel Gandy. A COVID-19 model for local authorities of the United Kingdom. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S1):S86–S95, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S86/7069480.

Gibson:2022:GJG

[415] Gavin J. Gibson. Gavin J. Gibson’s invited discussion: Contribution to the papers in session 2 of The Royal Statistical Society’s special topic meeting on Covid-19 transmission: 11 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S96–S98, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S96/7069469.

Nason:2022:GNI

[416] Guy Nason. Guy Nason’s invited discussion: Contribution to the papers in session 2 of The Royal Statistical Society’s special topic meeting on Covid-19 transmission: 11 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S99–S102, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S99/7069468.

Funk:2022:SFS

[417] Sebastian Funk, Sam Abbott, and Johannes Bracher. Sebastian Funk, Sam Abbott and Johannes Bracher’s discussion: Contribution to the papers in session 2 of The Royal Statistical Society’s special topic meeting on Covid-19 transmission: 11 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S103–S104, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S103/7069475.

Jewell:2022:CJD

[418] Christopher Jewell. Christopher Jewell’s discussion: Contribution to papers in session 2 of The Royal Statistical Society’s special topic meeting on COVID-19 transmission: 11 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S105–S106, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S105/7069484.

Teh:2022:ARD

[419] Yee Whye Teh, Bryn Elesedy, Bobby He, Michael Hutchinson, Sheheryar Zaidi, Avishkar Bhoopchand, Ulrich Paquet, Ne nad Tomasev, Jonathan Read, and Peter J. Diggle. Authors’ reply to the discussion of ‘Efficient Bayesian Inference of Instantaneous Reproduction Numbers at Fine Spatial Scales, with an Application to Mapping and Nowcasting the Covid-19 Epidemic in British Local Authorities’ by Teh et al. in session 2 of the Royal Statistical Society’s special topic meeting on COVID-19 transmission: 11 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S107–S109, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S107/7069492.

Mishra:2022:ARD

[420] Swapnil Mishra, James A. Scott, Daniel J. Laydon, Harrison Zhu, Neil M. Ferguson, Samir Bhatt, Seth Flaxman, and Axel Gandy. Authors’ reply to the discussion of ‘A COVID-19 Model for Local Authorities of the United Kingdom’ by Mishra et al. in session 2 of the Royal Statistical Society’s special topic meeting on COVID-19 transmission: 11 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S110–S111, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S110/7069499.

Pellis:2022:ERN

[421] Lorenzo Pellis, Paul J. Birrell, Joshua Blake, Christopher E. Overton, Francesca Scarabel, Helena B. Stage, Ellen Brooks-Pollock, Leon Danon, Ian Hall, Thomas A. House, Matt J. Keeling, Jonathan M. Read, JUNIPER Consortium, and Daniela De Angelis. Estimation of reproduction numbers in real time: Conceptual and statistical challenges. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S1):S112– S130, November 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/article/
185/Supplement_1/S112/7069502.

Dunbar:2022:AES

[422] Maria Bekker-Nielsen Dunbar, Felix Hofmann, Leonhard Held, and on Behalf of the Suspend modelling consortium. Assessing the effect of school closures on the spread of COVID-19 in Zurich. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S131–S142, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S131/7069501.

Kucharski:2022:AKI

[423] Adam Kucharski. Adam Kucharski’s invited discussion: Contribution to the papers in session 3 of The Royal Statistical Society’s special topic meeting on Covid-19 transmission: 11 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S143–S144, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        185/Supplement_1/S143/7069486.

Tomba:2022:GST

[424] Gianpaolo Scalia Tomba. Gianpaolo Scalia Tomba’s invited discussion: Contribution to the papers in session 3 of the Royal Statistical Society’s special topic meeting on Covid-19 transmission: 11 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S1):S145– S146, November 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

                                        academic.oup.com/jrsssa/article/
                                        185/Supplement_1/S145/7069500.

Diggle:2022:PJDc

[425] Peter J. Diggle. Peter J. Diggle’s discussion: Contribution to the papers in session 3 of the Royal Statistical Society’s special topic meeting on Covid-19 transmission: 11 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S1):S147, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        185/Supplement_1/S147/7069479.

Riley:2022:SRDb

[426] Steven Riley. Steven Riley’s discussion: Contribution to papers in session 3 of the Royal Statistical Society’s special topic meeting on COVID-19 transmission: 11 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S148–S149, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S148/7069474.

Panovska-Griffiths:2022:JPG

[427] Jasmina Panovska-Griffiths. Jasmina Panovska-Griffiths’ discussion contribution to papers in session 3 of The Royal Statistical Society’s special topic meeting on COVID-19 transmission: 11 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S1):S150– S151, November 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/article/
185/Supplement_1/S150/7069478.

Reynolds:2022:IRD

[428] Ian Reynolds. Ian Reynold’s discussion: Contribution to papers in session 3 of the Royal Statistical Society’s special topic meeting on COVID-19 transmission: 11 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S152, November 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_1/S152/7069483.

Pellis:2022:ARD

[429] Lorenzo Pellis, Paul J. Birrell, Joshua Blake, Ian Hall, Thomas A. House, Christopher E. Overton, Francesca Scarabel, Helena B. Stage, and Daniela De Angelis. Authors’ reply to the discussion of ‘Estimation of Reproduction Numbers in Real Time: Conceptual and Statistical Challenges’ by Pellis et al. in session 3 of The Royal Statistical Society’s special topic meeting on COVID-19 transmission: 11 June 2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S1):S153–S157, November 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        185/Supplement_1/S153/7069485.

Dunbar:2022:SRS

[430] Maria Bekker-Nielsen Dunbar, Felix Hofmann, and Leonhard Held. Session 3 of the RSS special topic meeting on Covid-19 transmission: Replies to the discussion. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S1):S158– S164, November 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

                                        academic.oup.com/jrsssa/article/
                                        185/Supplement_1/S158/7069490.

Anonymous:2022:IIf

[431] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S165–S166, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        185/Supplement_2/S165/7069505.

Jang:2022:BDM

[432] Don Jang and Ana Lucía Córdova Cazar. Big data meets survey science. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S167–S169, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S167/7069510.

Koebe:2022:IUU

[433] Till Koebe, Alejandra Arias-Salazar, Natalia Rojas-Perilla, and Timo Schmid. Intercensal updating using structurepreserving methods and satellite imagery. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S170–S196, December 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL

https://academic.oup.com/jrsssa/
article/185/Supplement_2/S170/7069525.

Consoli:2022:NFI

[434] Sergio Consoli, Luca Tiozzo Pezzoli, and Elisa Tosetti. Neural forecasting of the Italian sovereign bond market with economic news. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S197– S224, December 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/article/
185/Supplement_2/S197/7069532.

Keusch:2022:NPS

[435] Florian Keusch, Sebastian Bähr, Georg-Christoph Haas, Frauke Kreuter, Mark Trappmann, and Stephanie Eckman. Non-participation in smartphone data collection using research apps. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S225–S245, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S225/7069527.

Bach:2022:UPN

[436] Ruben L. Bach, Christoph Kern, Denis Bonnay, and Luc Kalaora. Understanding political news media consumption with digital trace data and natural language processing. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S246–S269, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        185/Supplement_2/S246/7069517.

Shook-Sa:2022:ENP

[437] Bonnie E. Shook-Sa, Michael G. Hudgens, Andrew L. Kavee, and David L. Rosen. Estimating the number of persons with HIV in jails via Web scraping and record linkage. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S270–S287, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        185/Supplement_2/S270/7069511.

Pezzoli:2022:SLH

[438] Luca Tiozzo Pezzoli and Elisa Tosetti. Seismonomics: Listening to the heartbeat of the economy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S288– S309, December 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

                                        academic.oup.com/jrsssa/article/
                                        185/Supplement_2/S288/7069546.

Kufner:2022:AES

[439] Benjamin Küfner, Joseph W. Sakshaug, and Stefan Zins. Analysing establishment survey non-response using administrative data and machine learning. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S310–S342, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S310/7069524.

Grow:2022:FAD

[440] André Grow, Daniela Perrotta, Emanuele Del Fava, Jorge Cimentada, Francesco Rampazzo, Sofia Gil-Clavel, Emilio Zagheni, René D. Flores, Ilana Ventura, and Ingmar Weber. Is Facebook’s advertising data accurate enough for use in social science research? Insights from a cross-national online survey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S343–S363, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S343/7069504.

Mneimneh:2022:ECL

[441] Zeina Mneimneh. Evaluation of consent to link Twitter data to survey data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S364–S386, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S364/7069515.

Silber:2022:LSD

[442] Henning Silber, Johannes Breuer, Christoph Beuthner, Tobias Gummer, Florian Keusch, Pascal Siegers, Sebastian Stier, and Bernd Weiß. Linking surveys and digital trace data: Insights from two studies on determinants of data sharing behaviour. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S387–S407, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        185/Supplement_2/S387/7069509.

Bosch:2022:WSS

[443] Oriol J. Bosch and Melanie Revilla. When survey science met Web tracking: Presenting an error framework for metered data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S408– S436, December 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

                                        academic.oup.com/jrsssa/article/
                                        185/Supplement_2/S408/7069520.

Ilic:2022:PIS

[444] Goran Ilic, Peter Lugtig, Barry Schouten, Maarten Streefkerk, Joris Mulder, Pradeep Kumar, and Seyit Höcük. Pictures instead of survey questions: an experimental investigation of the feasibility of using pictures in a housing survey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2): S437–S460, December 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        185/Supplement_2/S437/7069514.

Wu:2022:DFB

[445] Weichen Wu, Nynke Niezink, and Brian Junker. A diagnostic framework for the Bradley–Terry model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S461– S484, December 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/article/
185/Supplement_2/S461/7069516.

Gladwell:2022:PBE

[446] Daniel Gladwell, Gurleen Popli, and Aki Tsuchiya. Predictors of becoming not in education, employment or training: a dynamic comparison of the direct and indirect determinants. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S485–S514, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S485/7069508.

Lipowski:2022:FEI

[447] Cäcilia Lipowski, Ralf A. Wilke, and Bertrand Koebel. Fertility, economic incentives and individual heterogeneity: Register data-based evidence from France and Germany. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S515– S546, December 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/article/
185/Supplement_2/S515/7069543.

Rodriguez:2022:CCF

[448] Carlos E. Rodríguez and Ramsés H.

Mena. COVID-19 clinical footprint to infer about mortality. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S547– S572, December 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

                                        academic.oup.com/jrsssa/article/
                                        185/Supplement_2/S547/7069523.

Rudolph:2022:WED

[449] Kara E. Rudolph and Iván Díaz. When the ends do not justify the means: Learning who is predicted to have harmful indirect effects. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S573– S589, December 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

                                        academic.oup.com/jrsssa/article/
                                        185/Supplement_2/S573/7069507.

Dawber:2022:ERM

[450] James Dawber, Nicola Salvati, Enrico Fabrizi, and Nikos Tzavidis. Expectile regression for multi-category outcomes with application to small area estimation of labour force participation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S590–S619, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        185/Supplement_2/S590/7069506.

Cuellar:2022:PFC

[451] Maria Cuellar, Jacqueline Mauro, and Amanda Luby. A probabilistic formalisation of contextual bias: from forensic analysis to systemic bias in the criminal justice system. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S620–S643, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S620/7069513.

Yu:2022:BST

[452] Zhen Yu, Keming Yu, Wolfgang K.

Härdle, Xueliang Zhang, Kai Wang, and Maozai Tian. Bayesian spatiotemporal modeling for the inpatient hospital costs of alcohol-related disorders. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S644–S667, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S644/7069512.

Zhang:2022:SPA

[453] Bo Zhang and Eric J. Tchetgen Tchetgen. A semi-parametric approach to model-based sensitivity analysis in observational studies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S668– S691, December 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/article/
185/Supplement_2/S668/7069549.

Greene:2022:FKP

[454] Travis Greene, Galit Shmueli, Jan Fell, Ching-Fu Lin, and Han-Wei Liu. Forks over knives: Predictive inconsistency in criminal justice algorithmic risk assessment tools. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2): S692–S723, December 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        185/Supplement_2/S692/7069503.

Bikauskaite:2022:MMM

[455] Agne Bikauskaite, Isabel Molina, and Domingo Morales. Multivariate mixture model for small area estimation of poverty indicators. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2): S724–S755, December 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        185/Supplement_2/S724/7069526.

Simakani:2022:BRS

[456] June Elijah Simakani. Book review: Sampling Design and Analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S756, December 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        185/Supplement_2/S756/7069531.

Jen:2022:BRI

[457] Min-Hua Jen. Book review: Innovative Methods for Rare Disease Drug Development. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S757, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        185/Supplement_2/S757/7069542.

French:2022:BRP

[458] Simon French. Book review: Probability, Choice and Reason. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):

S758, December 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/article/
185/Supplement_2/S758/7069541.

Chowdhry:2022:BRH

[459] Amit K. Chowdhry. Book review: Handbook of Multiple Comparisons. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S759, December 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S759/7069540.

Pewsey:2022:BRO

[460] Arthur Pewsey. Book review: Object Oriented Data Analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S759–S760, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S759/7069530.

Glendinning:2022:BRR

[461] Richard H. Glendinning. Book review: Random Circulant Matrices. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S761–S762, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S761/7069539.

Chen:2022:BRR

[462] Li-Pang Chen. Book review: Replication and Evidence Factors in Observational Studies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S762– S763, December 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/article/
185/Supplement_2/S762/7069545.

Stein:2022:BRP

[463] Stefan Stein. Book review: Protecting Your Privacy in a Data-Driven World. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S763–S764, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S763/7069538.

Lazic:2022:BRB

[464] Stanley E. Lazic. Book review: Bayesian Modeling and Computation in Python Learning. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S764– S765, December 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/article/
185/Supplement_2/S764/7069537.

Greenacre:2022:BRR

[465] Michael Greenacre. Book review: R for Health Data Science. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S765– S766, December 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/article/
185/Supplement_2/S765/7069548.

Chen:2022:BRIb

[466] Li-Pang Chen. Book review: Introduction to High-Dimensional Statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S767, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S767/7069529.

McNicholl:2022:BRP

[467] Colin McNicholl. Book review: Project-Based R Companion to Introductory Statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S768, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S768/7069528.

Chen:2022:BRM

[468] Li-Pang Chen. Book review: Martingale Methods in Statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S769, December 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S769/7069547.

Todd:2022:BRS

[469] James Todd. Book review: Supervised Machine Learning for Text Analysis in R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S770, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S770/7069551.

Hossain:2022:BRL

[470] Md. Moyazzem Hossain. Book review: Linear Regression Models: Applications in R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S771, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        185/Supplement_2/S771/7069536.

Aalabaf-Sabaghi:2022:BRS

[471] Morteza Aalabaf-Sabaghi. Book review: Security Risk Models for Cyber Insurance. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S772, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        185/Supplement_2/S772/7069535.

Holmes:2022:BRS

[472] Dawn Holmes. Book review: Subjective Well-Being and Social Media. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S773, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        185/Supplement_2/S773/7069534.

Chowdhry:2022:BRM

[473] Amit K. Chowdhry. Book review: Machine Learning for Knowledge Discovery with R, Methodologies for Modeling, Inference and Prediction. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S774, December 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S774/7069533.

Aalabaf-Sabaghi:2022:BRH

[474] Morteza Aalabaf-Sabaghi. Book review: Handbook of Survival Analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S775, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S775/7069544.

Shalabh:2022:BRP

[475] Shalabh. Book review: Probability, Statistics, and Data: a Fresh Approach Using R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S776– S777, December 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/article/
185/Supplement_2/S776/7069519.

Chaturvedi:2022:BRH

[476] Anoop Chaturvedi. Book review: Handbook of Regression Analysis with Applications in R (Second Edition). Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S777–S778, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S777/7069518.

Lazic:2022:BRA

[477] Stanley E. Lazic. Book review: Artificial Intelligence and Causal Inference.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S778–S779, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S778/7069522.

Shalabh:2022:BRR

[478] Shalabh. Book review: Rank-Based Methods for Shrinkage and Selection: with Application to Machine Learning. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (S2):S779–S780, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
185/Supplement_2/S779/7069552.

Anonymous:2022:CVb

[479] Anonymous. Contents of volume 185, 2022. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S781–S789, December 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL

https://academic.oup.com/jrsssa/
article/185/Supplement_2/S781/7069550.

Jackson:2023:BMM

[480] Christopher Jackson, Belen Zapata-Diomedi, and James Woodcock. Bayesian multistate modelling of incomplete chronic disease burden data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186 (1):1–19, January 2023. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/article/
186/1/1/6998524.

Bennedsen:2023:MDS

[481] Mikkel Bennedsen, Eric Hillebrand, and Siem Jan Koopman. A multivariate dynamic statistical model of the global carbon budget 1959–2020. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(1): 20–42, January 2023. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/article/
186/1/20/7008652.

Hepler:2023:IAM

[482] Staci A. Hepler, David M. Kline, Andrea Bonny, Erin McKnight, and Lance A. Waller. An integrated abundance model for estimating county-level prevalence of opioid misuse in Ohio. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(1): 43–60, January 2023. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/article/
186/1/43/7010776.

Jeong:2023:MCC

[483] Himchan Jeong, George Tzougas, and Tsz Chai Fung. Multivariate claim count regression model with varying dispersion and dependence parameters. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(1):61–83, January 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/1/61/6999354.

Li:2023:MEM

[484] Hao Li, Daniel L. Millimet, and Punarjit Roychowdhury. Measuring economic mobility in India using noisy data: a partial identification approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(1):84–109, January 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/1/84/7008539.

Jenkins:2023:RRM

[485] Stephen P. Jenkins and Fernando Rios-Avila. Reconciling reports: modelling employment earnings and measurement errors using linked survey and administrative data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(1):110–136, January 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL

https://academic.oup.com/jrsssa/
article/186/1/110/7017750.

Egidi:2023:CSN

[486] Leonardo Egidi, Francesco Pauli, Nicola Torelli, and Susanna Zaccarin. Clustering spatial networks through latent mixture models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(1):137–156, January 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL

https://academic.oup.com/jrsssa/
article/186/1/137/6998435.

Smith:2023:DTH

[487] Fred Smith and Ian Diamond. David (Tim) Holt 1943–2022. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(1):157– 159, January 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL https://academic.
oup.com/jrsssa/article/186/1/157/

7008476. Updated by Ian Diamond (deceased 2019).

Coleman:2023:DHB

[488] Shirley Coleman. David Harold Baillie 1940–2021. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(1):159–160, January 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/1/159/6995399.

Bird:2023:ATL

[489] Sheila M. Bird and Vern T. Farewell.

Anthony (Tony) Leonard Johnson: founding editor of Statistics in Medicine and Medical Research Council biostatistician who helped to revolutionize the understanding and treatment of epilepsy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(1):160–162, January 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/1/160/6995401.

Jones:2023:PSM

[490] David Jones. Professor Sir Michael David Rawlins, 1941–2023. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(1):162– 163, January 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL https://academic.
oup.com/jrsssa/article/186/1/162/
7010778.

Dissanayake:2023:TSA

[491] Gnanadarsha Sanjaya Dissanayake.

Time-series analysis for the state-space model with R/Stan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(1):164–165, January 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic).  URL https://academic.
oup.com/jrsssa/article/186/1/164/
7017743.

Chowdhry:2023:MIM

[492] Amit K. Chowdhry. Multiple imputation of missing data in practice: Basic theory and analysis strategies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(1):165–166, January 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/1/165/7018801.

Chowdhry:2023:SMH

[493] Amit K. Chowdhry. Statistical methods for handling incomplete data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(1): 166, January 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL https://academic.
oup.com/jrsssa/article/186/1/166/
7016282.

Imai:2023:EEA

[494] Kosuke Imai, Zhichao Jiang, D. James Greiner, Ryan Halen, and Sooahn Shin. Experimental evaluation of algorithmassisted human decision-making: application to pretrial public safety assessment*. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2):167–189, April 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/2/167/7024674.

Cuellar:2023:PVT

[495] Maria Cuellar. Proposer of the vote of thanks and contribution to the discussion of Experimental evaluation of algorithm-assisted human decisionmaking: application to pretrial public safety assessment by Imai et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2):190–191, April 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/2/190/7036017.

Dawid:2023:SVT

[496] A. Philip Dawid. Seconder of the vote of thanks and contribution to the discussion of Experimental evaluation of algorithm-assisted human decisionmaking: application to pretrial public safety assessment by Imai et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2):192–193, April 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/2/192/7035413.

Python:2023:APC

[497] Andre Python. Andre Python’s contribution to the discussion of Experimental evaluation of algorithm-assisted human decision-making: application to pretrial public safety assessment by Imai et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2):194, April 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/2/194/7034887.

Ding:2023:PDC

[498] Peng Ding. Peng Ding’s contribution to the discussion of Experimental evaluation of algorithm-assisted human decision-making: application to pretrial public safety assessment by Imai et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2):195–198, April 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        186/2/195/7035414.

Conti:2023:PLC

[499] Pier Luigi Conti. Pier Luigi Conti’s contribution to the discussion of Experimental Evaluation of Algorithm-Assisted Human Decision-Making: Application to Pretrial Public Safety Assessment by Imai et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2):198– 199, April 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

                                        (electronic). URL https://academic.
                                        oup.com/jrsssa/article/186/2/198/
                                        7045508.

Sarvet:2023:ASM

[500] Aaron L. Sarvet and Mats J. Stensrud. Aaron Sarvet and Mats Stensrud’s contribution to the discussion of Experimental evaluation of algorithm-assisted human decision-making: application to pretrial public safety assessment by Imai et al., part 1. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2):199–200, April 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL

                                        https://academic.oup.com/jrsssa/
                                        article/186/2/199/7036026.

Stensrud:2023:MSV

[501] Mats J. Stensrud, Vanessa Didelez, and Aaron L. Sarvet. Mats Stensrud, Vanessa Didelez and Aaron Sarvet’s contribution to the discussion of Experimental evaluation of algorithm-assisted human decision-making: application to pretrial public safety assessment by Imai et al., part 2. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2):201–202, April 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL

https://academic.oup.com/jrsssa/
article/186/2/201/7036027.

Cortes:2023:ECC

[502] Efrén Cruz Cortés and Debashis Ghosh.

Efrén Cruz Cortés and Debashis Ghosh’s contribution to the discussion of Experimental evaluation of algorithm-assisted human-decision making: application to pretrial public safety assessment by Imai et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2):203–204, April 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/2/203/7039319.

Hunsicker:2023:EHC

[503] Eugenie Hunsicker. Eugenie Hunsicker’s contribution to the discussion of Experimental evaluation of algorithm-assisted human decision-making: application to pretrial public safety assessment by Imai et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2):204–205, April 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        186/2/204/7092819.

Hutton:2023:JLH

[504] Jane Luise Hutton. Jane Luise Hutton’s contribution to the discussion of Experimental evaluation of algorithm-assisted human decision-making: application to pretrial public safety assessment by Imai et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2):205–206, April 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        186/2/205/7036028.

Jose:2023:VRJ

[505] Victor Richmond R. Jose. Victor Richmond Jose’s contribution to the discussion of Experimental evaluation of algorithm-assisted human decisionmaking: application to pretrial public safety assessment by Imai et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2):206–207, April 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        186/2/206/7035415.

Kumar:2023:KKC

[506] Kuldeep Kumar. Kuldeep Kumar’s contribution to the discussion of Experimental evaluation of algorithm-assisted human-decision-making: application to pretrial public safety assessment by Imai et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2):208, April 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/2/208/7049148.

VanderWeele:2023:TVC

[507] Tyler J. VanderWeele. Tyler Vander-Weele’s contribution to the discussion of Experimental evaluation of algorithmassisted human decision-making: application to pretrial public safety assessment by Imai et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2):209–210, April 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL

https://academic.oup.com/jrsssa/
article/186/2/209/7035417.

Wijayatunga:2023:PWC

[508] Priyantha Wijayatunga. Priyantha Wijayatunga’s contribution to the discussion of Experimental evaluation of algorithm-assisted human decisionmaking: application to the pretrial safety assessment by Imai et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2): 210–211, April 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL https://academic.
oup.com/jrsssa/article/186/2/210/
7036029.

Imai:2023:ARD

[509] Kosuke Imai, Zhichao Jiang, D. James Greiner, Ryan Halen, and Sooahn Shin. Authors’ reply to the discussion of Experimental evaluation of algorithmassisted human decision-making: application to pretrial public safety assessment. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2):212–216, April 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        186/2/212/7056267.

Permanyer:2023:GIB

[510] Iñaki Permanyer, Isaac Sasson, and Francisco Villavicencio. Group- and individual-based approaches to health inequality: towards an integration. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2):217–240, April 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        186/2/217/7023973.

Advani:2023:MTI

[511] Arun Advani, Andy Summers, and Hannah Tarrant. Measuring top income shares in the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2):241–258, April 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL

                                        https://academic.oup.com/jrsssa/
                                        article/186/2/241/7045959.

Hewson:2023:TUI

[512] Paul Hewson. Targeting uplift: an Introduction to Net Scores by R. Michel, I. Schnakenburg, and T. von Martens. Cham, Switzerland, Springer, 2019. pp.

384. ISBN-13: 978-3-030-22627-5. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(2): 259, April 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

                                        (electronic). URL https://academic.
                                        oup.com/jrsssa/article/186/2/259/
                                        7030925.

Vinciotti:2023:NS

[513] Veronica Vinciotti, Luigi Augugliaro, and Ernst Wit. Networks and society. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(3):261–263, July 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/3/261/7160245.

DeBenedictis:2023:CNC

[514] Luca De Benedictis, Roberto Rondinelli, and Veronica Vinciotti. Cultures as networks of cultural traits: a unifying framework for measuring culture and cultural distances. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(3):264–293, July 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL

https://academic.oup.com/jrsssa/
article/186/3/264/7043260.

Shenvi:2023:BDS

[515] Aditi Shenvi, Francis Oliver Bunnin, and Jim Q. Smith. A Bayesian decision support system for counteracting activities of terrorist groups. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(3): 294–312, July 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL https://academic.
oup.com/jrsssa/article/186/3/294/
6999023.

Gilardi:2023:MEM

[516] Andrea Gilardi, Riccardo Borgoni, Luca Presicce, and Jorge Mateu. Measurement error models for spatial network lattice data: Analysis of car crashes in Leeds. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(3):313–334, July 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/3/313/7146735.

Dey:2023:DTT

[517] Asim K. Dey, Stephen J. Young, and Yulia R. Gel. From Delaunay triangulation to topological data analysis: generation of more realistic synthetic power grid networks. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(3):335–354, July 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/3/335/7175407.

DeBacco:2023:LNM

[518] Caterina De Bacco, Martina Contisciani, Jonathan Cardoso-Silva, Hadiseh Safdari, Gabriela Lima Borges, Diego Baptista, Tracy Sweet, Jean-Gabriel Young, Jeremy Koster, Cody T. Ross, Richard McElreath, Daniel Redhead, and Eleanor A. Power. Latent network models to account for noisy, multiply reported social network data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(3): 355–375, July 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL https://academic.
oup.com/jrsssa/article/186/3/355/
7030939.

Koskinen:2023:MLA

[519] Johan Koskinen and Tom A. B. Snijders. Multilevel longitudinal analysis of social networks. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(3):376–400, July 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/3/376/6998523.

Giudici:2023:NSE

[520] Paolo Giudici, Paolo Pagnottoni, and Alessandro Spelta. Network self-exciting point processes to measure health impacts of COVID-19. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(3):401–421, July 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL

https://academic.oup.com/jrsssa/
article/186/3/401/7008583.

Bartolucci:2023:TAH

[521] Francesco Bartolucci, Paolo Li Donni, and Antonietta Mira. Temporal analysis of hospital network data by hierarchical Bayesian p_{2} models with covariates. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(3):422–440, July 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/3/422/7096467.

Suveges:2023:NCE

[522] Maria Süveges and Sofia Charlotta Olhede. Networks with correlated edge processes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(3):441–462, July 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/3/441/7083542.

Zhu:2023:DPN

[523] Xiaojing Zhu, Cantay Caliskan, Dino P. Christenson, Konstantinos Spiliopoulos, Dylan Walker, and Eric D. Kolaczyk. Disentangling positive and negative partisanship in social media interactions using a coevolving latent space network with attractors model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(3):463–480, July 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        186/3/463/7058002.

DAngelo:2023:MBC

[524] Silvia D’Angelo, Marco Alfò, and Michael Fop. Model-based clustering for multidimensional social networks. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(3):481–507, July 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        186/3/481/6998598.

Artico:2023:DLS

[525] I. Artico and E. C. Wit. Dynamic latent space relational event model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(3):508–529, July 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

                                        /academic.oup.com/jrsssa/article/
                                        186/3/508/7126053.

Basini:2023:ACB

[526] Francesca Basini, Vasiliki Tsouli, Ioannis Ntzoufras, and Nial Friel. Assessing competitive balance in the English Premier League for over forty seasons using a stochastic block model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(3):530–556, July 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/3/530/7035692.

Fritz:2023:MLD

[527] Cornelius Fritz, Giacomo De Nicola, Sevag Kevork, Dietmar Harhoff, and Göran Kauermann. Modelling the large and dynamically growing bipartite network of German patents and inventors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(3):557–576, July 2023. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https:/

/academic.oup.com/jrsssa/article/
186/3/557/7071940.

Lerner:2023:RHM

[528] Jürgen Lerner and Alessandro Lomi. Relational hyperevent models for polyadic interaction networks. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(3):577–600, July 2023. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic).  URL https://academic.
oup.com/jrsssa/article/186/3/577/
6998599.

Diamond:2024:FES

[529] Ian Diamond, Grant Fitzner, Richard Heys, Michael Keoghan, and Darren Morgan. The future of economic statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):1–2, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

                                        /academic.oup.com/jrsssa/article/
                                        187/1/1/7186556.

Dunne:2024:SPE

[530] John Dunne and Li-Chun Zhang. A system of population estimates compiled from administrative data only. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):3–21, January 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

                                        /academic.oup.com/jrsssa/article/
                                        187/1/3/7142865.

Abbott:2024:PVT

[531] Owen Abbott. Proposer of the vote of thanks to Dunne and Zhang and contribution to the discussion of ‘A system of population estimates compiled from administrative data only’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):22–23, January 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

                                        /academic.oup.com/jrsssa/article/
                                        187/1/22/7238223.

Shlomo:2024:SVT

[532] Natalie Shlomo. Seconder of the vote of thanks to Dunne and Zhang and contribution to the discussion of ‘A system of population estimates compiled from administrative data only’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):23–25, January 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

                                        /academic.oup.com/jrsssa/article/
                                        187/1/23/7233304.

Baffour:2024:BBC

[533] Bernard Baffour. Bernard Baffour’s contribution to the discussion of ‘A system of population estimates compiled from administrative data only’ by Dunne and Zhang. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):25–26, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/1/25/7234284.

Chind:2024:DAC

[534] Arun Chind. Dr Arun Chind’s contribution to the discussion of ‘A system of population estimates compiled from administrative data only’ by Dunne and Zhang. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):26–27, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/1/26/7257479.

Smith:2024:PSC

[535] Paul A. Smith. Paul Smith’s contribution to the discussion of ‘A system of population estimates compiled from administrative data only’ by Dunne and Zhang. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):27–28, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/1/27/7233303.

Bernardini:2024:ABN

[536] Antonella Bernardini, Nicoletta Cibella, and Fabrizio Solari. Antonella Bernadini, Nicoletta Cibella, and Fabrizio Solari’s contribution to the discussion of ‘A system of population estimates compiled from administrative data only’ by Dunne and Zhang. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):28–29, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/1/28/7235932.

Law:2024:ELC

[537] Eleanor C. Law. Eleanor Law’s contribution to the discussion of ‘A system of population estimates compiled from administrative data only’ by Dunne and Zhang. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):29–30, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/1/29/7234931.

Allin:2024:PAC

[538] Paul Allin. Paul Allin’s contribution to the discussion of ‘A system of population estimates compiled from administrative data only’ by Dunne and Zhang. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):30, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/1/30/7237415.

Chai:2024:CPC

[539] Christine P. Chai. Christine P. Chai’s contribution to the discussion of “A system of population estimates compiled from administrative data only” by Dunne and Zhang. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):31– 32, January 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL http://academic.
oup.com/jrsssa/article/187/1/31/
7234932.

Henry:2024:SHK

[540] Sarah Henry and Katie O’Farrell. Sarah Henry and Katie O’Farrell’s contribution to the discussion of ‘A system of population estimates compiled from administrative data only’ by Dunne and Zhang. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):32–33, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/1/32/7241514.

Pfeffermann:2024:DPC

[541] Danny Pfeffermann. Danny Pfeffermann’s contribution to the discussion of ‘A system of population estimates compiled from administrative data only’ by Dunne and Zhang. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):33– 34, January 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL http://academic.
oup.com/jrsssa/article/187/1/33/
7233302.

Woolford:2024:JWC

[542] Jennet Woolford. Jennet Woolford’s contribution to the discussion of ‘A system of population estimates compiled from administrative data only’ by Dunne and Zhang. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):34–35, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

                                        /academic.oup.com/jrsssa/article/
                                        187/1/34/7232259.

Dunne:2024:JDL

[543] John Dunne and Li-Chun Zhang. John Dunne and Li-Chun Zhang’s reply to the discussion of ‘A system of population estimates compiled from administrative data only’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):35–38, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

                                        /academic.oup.com/jrsssa/article/
                                        187/1/35/7274076.

Mareeswaran:2024:NMS

[544] M. Mareeswaran, Shubhajit Sen, and Soudeep Deb. New methods of structural break detection and an ensemble approach to analyse exchange rate volatility of Indian rupee during coronavirus pandemic. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):39– 61, January 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

                                        (electronic). URL http://academic.
                                        oup.com/jrsssa/article/187/1/39/
                                        7197368.

Guha:2024:SAE

[545] Saurav Guha and Hukum Chandra. Small area estimation under a spatially correlated multivariate area-level model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):62–84, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

                                        /academic.oup.com/jrsssa/article/
                                        187/1/62/7202186.

DeNicolo:2024:SAE

[546] Silvia De Nicolò, Maria Rosaria Ferrante, and Silvia Pacei. Small area estimation of inequality measures using mixtures of Beta. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):85– 109, January 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL http://academic.
oup.com/jrsssa/article/187/1/85/
7227743.

Bhuyan:2024:EPS

[547] Prajamitra Bhuyan and Kiranmoy Chatterjee. Estimation of population size with heterogeneous catchability and behavioural dependence: applications to air- and water-borne disease surveillance. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):110–131, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/1/110/7231027.

Armillotta:2024:UVH

[548] Mirko Armillotta, Konstantinos Fokianos, and Andrea Guizzardi. Unveiling Venice’s hotels competition networks from dynamic pricing digital market. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):132–159, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/1/132/7233171.

Gao:2024:NPP

[549] Jiti Gao, Oliver Linton, and Bin Peng.

A non-parametric panel model for climate data with seasonal and spatial variation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):160–179, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/1/160/7236265.

Spelta:2024:WBL

[550] Alessandro Spelta and Nicolò Pecora. Wasserstein barycenter for link prediction in temporal networks. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):180– 208, January 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL http://academic.
oup.com/jrsssa/article/187/1/180/
7235774.

Bach:2024:HUG

[551] Philipp Bach, Victor Chernozhukov, and Martin Spindler. Heterogeneity in the US gender wage gap. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187 (1):209–230, January 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/1/209/7239118.

Carrino:2024:MSI

[552] Ludovico Carrino, Luca Farnia, and Silvio Giove. Measuring social inclusion in Europe: a non-additive approach with the expert-preferences of public policy planners. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):231–259, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/1/231/7260716.

Kumar:2024:SA

[553] Kuldeep Kumar. Spatial analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1): 260, January 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL http://academic.
oup.com/jrsssa/article/187/1/260/
7152366.

Kumar:2024:DDL

[554] Kuldeep Kumar. Dicing with death living by data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):260–261, January 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic).  URL http://academic.
oup.com/jrsssa/article/187/1/260/
7152367.

Aalabaf-Sabaghi:2024:RAT

[555] Morteza Aalabaf-Sabaghi. Risk assessment: Theory, methods, and applications. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):261–262, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/1/261/7205356.

Kalyani:2024:BRR

[556] V. Kalyani. Book review: Regression Analysis in R: a Comprehensive View for the Social Sciences, by Jocelyn H. Bolin, 1st ed., CRC Press, 2023. xii + 180 pp, ISBN 978-1-032-25775-4, £145.00 (Hardback), £51.29 (eBook). Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):262–263, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/1/262/7219780.

Glendinning:2024:GFD

[557] Richard H. Glendinning. Geostatistical functional data analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):263– 264, January 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL http://academic.
oup.com/jrsssa/article/187/1/263/
7219779.

Thangam:2024:PAU

[558] Dhanabalan Thangam. Predictive analytics using statistics and big data: Concepts and modelling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):264– 265, January 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL http://academic.
oup.com/jrsssa/article/187/1/264/
7235725.

Lazic:2024:BRS

[559] Stanley E. Lazic. Book review: Statistical Methods for Mediation, Confounding and Moderation Analysis Using R and SAS, by Qingzhao Yu and Bin Li, CRC Press, 2022. 294pp, ISBN: 978-0-367-36547-9, £125.00. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1): 265, January 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL http://academic.
oup.com/jrsssa/article/187/1/265/
7235727.

Kalyani:2024:BRC

[560] V. Kalyani. Book review: Crime by the Numbers: a Criminologist’s Guide to R, by Jacob Kaplan, 1st Edition, 2023, Chapman and Hall/CRC, xxi + 432 pp, £59.99 (Hardback) and £59.99 (eBook), ISBN 978-1-003-27921-1. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):266, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/1/266/7235728.

Srivastava:2024:BD

[561] Mukesh Srivastava. Biostatistics decoded. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):266–267, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/1/266/7239119.

Hutchison:2024:HSM

[562] Dougal Hutchison. Handbook of statistical methods for randomized controlled trials. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):267–268, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/1/267/7246316.

Anonymous:2024:CRC

[563] Anonymous. Calyampudi Radhakrishna (CR) Rao 1920–2023. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):269– 271, January 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

                                        (electronic). URL http://academic.
                                        oup.com/jrsssa/article/187/1/269/
                                        7407358.

Anonymous:2024:R

[564] Anonymous. Referees. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):272– 278, January 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

                                        (electronic). URL http://academic.
                                        oup.com/jrsssa/article/187/1/272/
                                        7516313.

Steele:2024:LAE

[565] Fiona Steele, Siliang Zhang, Emily Grundy, and Tania Burchardt. Longitudinal analysis of exchanges of support between parents and children in the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):279–304, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

                                        /academic.oup.com/jrsssa/article/
                                        187/2/279/7248483.

Nakamura:2024:DEI

[566] Ryota Nakamura, Andrea Albanese, Emma Coombes, and Marc Suhrcke. Do economic incentives promote physical activity? Evidence from the London Congestion Charge. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):305– 320, April 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

                                        (electronic). URL http://academic.
                                        oup.com/jrsssa/article/187/2/305/
                                        7246822.

Lan:2024:MSM

[567] Zhou Lan and Le Bao. Multivariate spatial modelling for predicting missing HIV prevalence rates among key populations. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):321–337, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/2/321/7263079.

Pillinger:2024:DSR

[568] Rebecca Pillinger, Fiona Steele, George Leckie, and Jennifer Jenkins. A dynamic social relations model for clustered longitudinal dyadic data with continuous or ordinal responses. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):338– 357, April 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL http://academic.
oup.com/jrsssa/article/187/2/338/
7260780.

Wang:2024:RPR

[569] Lingxiao Wang, Yan Li, Barry I.

Graubard, and Hormuzd A. Katki. Representative pure risk estimation by using data from epidemiologic studies, surveys, and registries: estimating risks for minority subgroups. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):358– 373, April 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL http://academic.
oup.com/jrsssa/article/187/2/358/
7405297.

Park:2024:GAI

[570] Chan Park and Hyunseung Kang. A groupwise approach for inferring heterogeneous treatment effects in causal inference. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):374–392, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

                                        /academic.oup.com/jrsssa/article/
                                        187/2/374/7335250.

Barons:2024:WBS

[571] Martine J. Barons and Aditi Shenvi. Where the bee sucks: a dynamic Bayesian network approach to decision support for pollinator abundance strategies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):393–409, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

                                        /academic.oup.com/jrsssa/article/
                                        187/2/393/7336929.

Castro:2024:BAE

[572] Andrea Aparicio Castro, Arkadiusz Wiśniowski, and Francisco Rowe. A Bayesian approach to estimate annual bilateral migration flows for South America using census data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):410– 435, April 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

                                        (electronic). URL http://academic.
                                        oup.com/jrsssa/article/187/2/410/
                                        7420535.

Goldstein:2024:ITV

[573] Isaac H. Goldstein, Jon Wakefield, and Volodymyr M. Minin. Incorporating testing volume into estimation of effective reproduction number dynamics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):436–453, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/2/436/7471058.

Bregantini:2024:BCP

[574] Daniele Bregantini, Laetitia H. M.

Schmitt, and Jacco J. J. Thijssen. A Bayesian change-point detection approach to the economic evaluation of risky projects: an application to healthcare technology assessment. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):454– 476, April 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL http://academic.
oup.com/jrsssa/article/187/2/454/
7442048.

Koop:2024:ISD

[575] Gary Koop, Stuart McIntyre, James Mitchell, Aubrey Poon, and Ping Wu. Incorporating short data into large mixed-frequency vector autoregressions for regional nowcasting. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):477– 495, April 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL http://academic.
oup.com/jrsssa/article/187/2/477/
7441643.

Stephenson:2024:IDC

[576] Briana J. K. Stephenson, Stephanie M.

Wu, and Francesca Dominici. Identifying dietary consumption patterns from survey data: a Bayesian nonparametric latent class model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2): 496–512, April 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL http://academic.

                                        oup.com/jrsssa/article/187/2/496/
                                        7471024.

Divekar:2024:RTF

[577] Chinmay Divekar, Soudeep Deb, and Rishideep Roy. Real-time forecasting within soccer matches through a Bayesian lens. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):513–540, April 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

                                        tronic).  URL http://academic.
                                        oup.com/jrsssa/article/187/2/513/
                                        7512935.

McCulloch:2024:MMP

[578] Andrew McCulloch. Measurement models for psychological attributes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2): 541, April 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

                                        (electronic). URL http://academic.
                                        oup.com/jrsssa/article/187/2/541/
                                        7247587.

McCulloch:2024:PSS

[579] Andrew McCulloch. The psychometrics of standard setting. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):542, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

                                        /academic.oup.com/jrsssa/article/
                                        187/2/542/7247586.

Kalyani:2024:BDS

[580] V. Kalyani. Big data and social science data science methods and tools for research and practice. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):542– 543, April 2024. CODEN JSSAEF.

ISSN 0964-1998 (print), 1467-985X

(electronic). URL http://academic.
oup.com/jrsssa/article/187/2/542/
7242403.

Reese:2024:DSE

[581] R. Allan Reese. Data science ethics:

Concepts, techniques and cautionary tales. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):543–544, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/2/543/7242405.

Pinon:2024:BRH

[582] Adelson Piñón. Book review: Handbook of Regression Modeling in People Analytics, With Examples in R and Python, by Keith McNulty, 1st edition, CRC Press. 2021. 255 pp. ISBN: 978-1-032-04174-2, £59.99 (hbk). Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):544– 545, April 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL http://academic.
oup.com/jrsssa/article/187/2/544/
7256154.

Aalabaf-Sabaghi:2024:BRL

[583] Morteza Aalabaf-Sabaghi. Book review: Learning Microeconometrics with R, by Christopher P Adams , CRC Press , Chapman & Hall. 2021. xxx + 368pp. ISBN: 978-0-367-25538-1, £77.99 (hbk). Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):545–546, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/2/545/7263080.

Pinon:2024:NSI

[584] Adelson Piñón. Number savvy: From the invention of numbers to the future of data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):546–547, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

                                        /academic.oup.com/jrsssa/article/
                                        187/2/546/7256167.

Murugaiah:2024:PSE

[585] Suryakumar Murugaiah. Perspectives in sustainable equity investing. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2): 547–548, April 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

                                        (electronic). URL http://academic.
                                        oup.com/jrsssa/article/187/2/547/
                                        7283028.

Shalabh:2024:AUB

[586] Shalabh. Applied univariate, bivariate, and multivariate statistics using Python: a Beginner’s guide to advanced data analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):548, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

                                        /academic.oup.com/jrsssa/article/
                                        187/2/548/7282112.

Chaturvedi:2024:ISD

[587] Anoop Chaturvedi. Introduction to statistics and data analysis (with exercises, solutions and applications in r). Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):549, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/2/549/7321171.

Dietz:2024:RMI

[588] Sebastian Dietz. Risk measures and insurance solvency benchmarks. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2): 549–550, April 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X

(electronic). URL http://academic.
oup.com/jrsssa/article/187/2/549/
7424804.

Lazic:2024:BRE

[589] Stanley E. Lazic. Book review: Exploring Modeling with Data and Differential Equations Using R, by John Zobitz, 2022, Chapman & Hall/CRC Press, Boca Raton, FL, 356 pages, £74.99 (Hardcover), ISBN: 978-1-032-25948-2. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):550–551, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/2/550/7420307.

Chaturvedi:2024:TSD

[590] Anoop Chaturvedi. Time series for data sciences: Analysis and forecasting. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):551–552, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/2/551/7442025.

Dalal:2024:CLM

[591] Siddhartha Dalal and James Landwehr.

Colin Lingwood Mallows: a great ambidextrous statistician, 1930–2023. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):553–554, April 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

                                        /academic.oup.com/jrsssa/article/
                                        187/2/553/7486387.

Anonymous:2024:JMB

[592] Anonymous. John Morgan Bynner 1938–2023. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):555, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

                                        /academic.oup.com/jrsssa/article/
                                        187/2/555/7492510.

Muller:2024:TG

[593] Hans-Georg Müller. Theo Gasser, 1941– 2023. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):556–557, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

                                        /academic.oup.com/jrsssa/article/
                                        187/2/556/7513468.

Smith:2024:LB

[594] Paul A. Smith. Lawrence S. Berman 1928–2023. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):557–558, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

                                        /academic.oup.com/jrsssa/article/
                                        187/2/557/7585210.

Obe:2024:RM

[595] Geoffrey C. Reed Obe. Rita Maurice 1929–2023. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):558–560, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/2/558/7608749.

Gavras:2024:CIC

[596] Konstantin Gavras, Jan Karem Höhne, Annelies G. Blom, and Harald Schoen. Correction to: Innovating the collection of open-ended answers: The linguistic and content characteristics of written and oral answers to political attitude questions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):561–562, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:/

/academic.oup.com/jrsssa/article/
187/2/561/7624468. See [297].